### v1

# Understanding New-Knowledge-Induced Factual Hallucinations in LLMs: Analysis, Solution and Interpretation

Abstract
Previous studies show that introducing new knowledge during large language models (LLMs) fine-tuning can lead to the generation of erroneous output when tested on known information, thereby triggering factual hallucinations. However, existing studies have not deeply investigated the specific manifestations and underlying mechanisms of these hallucinations. Our work addresses this gap by designing a controlled dataset Biography-Reasoning, and conducting a fine-grained analysis across multiple knowledge types and two task types, including knowledge question answering (QA) and knowledge reasoning tasks. We find that when fine-tuned on a dataset in which a specific knowledge type consists entirely of new knowledge, LLMs exhibit significantly increased hallucination tendencies. This suggests that the high unfamiliarity of a particular knowledge type, rather than the overall proportion of new knowledge, is a stronger driver of hallucinations, and these tendencies can even affect other knowledge types in QA tasks. To mitigate such factual hallucinations, we propose KnownPatch, which patches a small number of known knowledge samples in the later stages of training, effectively alleviating new-knowledge-induced hallucinations. Through attention analysis, we find that learning new knowledge reduces the model’s attention to key entities in the question, thus causing excessive focus on the surrounding context, which may increase the risk of hallucination. Moreover, the attention pattern can propagate to similar contexts, facilitating the spread of hallucinations to textually similar questions. Our method effectively mitigates the disruption of new knowledge learning to the model’s attention on key entities, accompanied by improved performance.
Codehttps://github.com/NJUNLP/New-Knowledge-Induced-Factual-Hallucinations
1 Introduction
LLMs embed rich factual knowledge in their parameters during pre-training on massive text corpora (Petroni et al. 2019; Cohen et al. 2023). Subsequently, post-training enables them to learn to follow human instructions and exhibit superior performance across various downstream tasks (Ouyang et al. 2022; Wei et al. 2022).
Figure 1: The impact of learning new knowledge on attention patterns and hallucination behavior. When an model is trained on unknown facts, it may be more prone to produce factual hallucinations, the severity of which correlates with the degree of attention paid to key entities and is modulated by contextual similarity. By injecting a small amount of known knowledge at the end of training via KnowPatch, this issue can be effectively mitigated.
However, during the Supervised Fine-Tuning (SFT) phase, models may encounter new knowledge not covered in pre-training. Prior researches (Ghosal, Hashimoto, and Raghunathan 2024; Lin et al. 2023; Ovadia et al. 2023; Gekhman et al. 2024; Sun et al. 2025) suggest that introducing new knowledge in the post-training phase increases the risk of factual hallucinations, where models generate fabricated yet plausible statements. This occurs because, when models learn new facts absent from pre-training, they may erroneously generate related information in irrelevant contexts (Gekhman et al. 2024; Sun et al. 2025). These studies primarily focus on the effects within knowledge-intensive QA tasks during SFT, and we advance this line of research by conducting a systematic analysis of fine-grained manifestations and underlying causes of hallucinations.
To enable a comprehensive investigation, we construct a controlled experimental dataset, Biography-Reasoning. The dataset is composed of biographical entities and their four attributes (birth, death, major, and university), which serve as four knowledge types. We further design twelve reasoning tasks using these knowledge. By controlling the proportion of known and unknown knowledge within different types and tasks in the training data, we systematically investigate the impact of learning new knowledge on hallucination risks.
Our experiments reveal that, for knowledge QA tasks, training on unknown knowledge significantly elevates hallucination risks in the same type test set, with some cross-type influence on other QA test sets; in tasks involving knowledge reasoning, training on reasoning tasks containing unknown knowledge primarily affects the same reasoning task, with some impact on knowledge QA and limited influence on other reasoning tasks. Importantly, we find that if a knowledge type consists entirely of new knowledge, even a small amount of such data can markedly increase hallucination tendencies. This poses challenges for non-perfect filtering-based hallucination mitigation methods: inadvertently retained small amounts of unknown data can trigger severe hallucinations.
To address this challenge, we propose KnownPatch, a simple yet effective strategy that places a small number of known knowledge in later stages of training. Instead of relying on exhaustive filtering of unknown content, our method stabilizes learning by reinducing known knowledge late in training. Experiments show that even a small injection of known data significantly reduces hallucination tendencies across diverse settings.
Finally, we perform an interpretability analysis of the model’s attention distribution. The results reveal that learning new knowledge significantly weakens the model’s attention to key entities in the question, instead may erroneously bind new knowledge to other contexts, thereby triggering factual hallucinations. Therefore, tasks that share similar contexts are more likely to be affected and exhibit stronger hallucination tendencies. KnownPatch effectively restores and enhances the model’s attention to key entities, thereby mitigating hallucination risks. These findings are visually presented in Figure 1.
In summary, the main contributions of this paper are:
  • Fine-Grained Analysis: A detailed analysis across knowledge types and task types reveals the manifestations of new-knowledge-induced hallucinations, showing that when all knowledge within a specific type is entirely unknown, it is more likely to trigger severe hallucinations, even on unrelated QA test sets.
  • Mitigation Strategy: The introduction of KnownPatch, a training method that places a small amount of known knowledge in the later stages, significantly reducing hallucination risks without data modifications.
  • Mechanism Interpretability: An analysis of attention mechanisms shows that learning new knowledge reduces attention to key question entities and erroneously binds new knowledge to contexts, causing hallucinations. Moreover, similar contexts facilitate the spread of these altered attention patterns, enabling cross-task hallucination effects.
2 Related Work
New Knowledge and Hallucinations
Existing studies have indicated that introducing new knowledge into LLMs may trigger hallucinations (Ghosal, Hashimoto, and Raghunathan 2024; Lin et al. 2023; Ovadia et al. 2023). Subsequent works have provided deeper analyses of this phenomenon(Gekhman et al. 2024; Kang et al. 2024; Sun et al. 2025). Gekhman et al. (2024) found that as the proportion of new knowledge in fine-tuning data increases, the model’s hallucination tendency intensifies. Kang et al. (2024) analyzed that when fine-tuned LLMs encounter unknown queries during testing, their responses imitate those associated with unknown examples in the fine-tuning data. Sun et al. (2025), from the perspective of token probabilities, examined how, after learning new knowledge, the generation probability of relevant answer entity tokens in irrelevant contexts increases significantly, suggesting that the model may erroneously generalize knowledge, leading to hallucinations. However, previous studies focus mainly on closed-book QA settings with mixed knowledge types during training, while our controlled setup disentangles them to provide a more detailed analysis of new knowledge-induced hallucinations across types and tasks. Furthermore, we also investigate the underlying mechanisms of these phenomenon through an analysis of attention weights.
Reducing Hallucinations
Numerous studies are currently exploring ways to mitigate model hallucinations. A common approach involves providing additional relevant context to the model to reduce hallucinations during generation, such as through retrieval from knowledge bases or leveraging other large models to generate context(Shuster et al. 2021; Sun et al. 2022; Asai et al. 2024; Feng et al. 2023). Additionally, some research explicitly avoids hallucination risks by refusing to answer uncertain or unfamiliar questions(Yadkori et al. 2024; Zhu et al. 2025; Duwal 2025). In another direction, many studies encourage the model to generate more known knowledge from pre-training, for example, by promoting factual outputs via reinforcement learning (Rafailov et al. 2023; Kang et al. 2024; Li and Ng 2025; Gu et al. 2025) or by training only on known knowledge during supervised fine-tuning (Lin et al. 2024; Ghosal, Hashimoto, and Raghunathan 2024; Liu et al. 2024) to enhance the model. Our work builds on SFT with known knowledge approach, but rather than pursuing comprehensive filtering across all training data, KnownPatch only introduces a small number of known knowledge samples in the later stages of training, and alleviates the model’s tendency for hallucination.
3 Methodology of Analyzing Hallucinations
We aim to systematically investigate factual hallucinations in LLMs caused by learning different knowledge-related tasks. However, in real-world datasets, most factual knowledge may have already been seen by LLMs during pre-training, making it difficult to precisely control whether the knowledge being learned is new to the model. To address this limitation, we construct a controlled experimental environment and a synthetic dataset named Biography-Reasoning, which allows a controllable examination of hallucination behaviors under varying knowledge types and task types.
3.1 Biography-Reasoning Dataset
Following the data construction methodologies of Allen-Zhu and Li (2024); Zheng et al. (2025), we design the Biography-Reasoning dataset. The dataset centers on individuals as the key entities, with each person associated with four attributes: birth year (B), death year (D), major (M), and university (U). We refer to the same attribute of different individuals as a knowledge type.
Our dataset includes two main types of knowledge-related tasks, i.e. knowledge QA and knowledge-based reasoning tasks. For knowledge QA tasks, questions are formulated by directly querying one of the attributes given the person’s name. Each task consists of questions on a single type, resulting in four QA tasks (B_QA, D_QA, M_QA and U_QA).
For knowledge-based reasoning tasks, we design three fundamental types of chain-of-thought-requiring reasoning tasks. Specifically, these include:
  • SR (Single Reasoning): extracting one attribute from a single entity and performing a simple reasoning process;
  • CR (Comparative Reasoning): extracting one attribute from each of two entities and performing comparative reasoning between them;
  • NR (Novel Reasoning): extracting one attribute from a single entity and performing a newly defined reasoning task, such as mathematical or symbolic reasoning.
Table 1 presents examples of the constructed questions. Some of the reasoning tasks are intentionally designed to be more complex than mere knowledge extraction as QA problems. They require further reasoning, as well as auxiliary knowledge (e.g., the Major Dentistry belongs to the field Medicine), which the model is expected to contain. To further guarantee the model’s proficiency, we additionally collect and train on these auxiliary facts.
For each knowledge type we construct one QA and three reasoning tasks, leading to a total of four QA and 12 reasoning tasks per individual. Further dataset details can be found in Appendix A.
Category Example
M_QA Question: What major did Darreus Hsiao study?
Answer: Dentistry
M_SR Question: What field does Darreus Hsiao’s major belong to?
Answer: Darreus Hsiao’s major is Dentistry. Dentistry belongs to Medicine.
The answer is: Medicine
M_CR Question: Do Darreus Hsiao and Virgus Hong’s majors belong to the same field?
Answer: Darreus Hsiao’s major is Dentistry. Dentistry belongs to Medicine. Virgus Hong’s major is Nursing. Nursing belongs to Medicine. Medicine and Medicine are the same.
The answer is: YES
M_NR Question: What is the sequence of odd-positioned letters in the first word of Darreus Hsiao’s major name?
Answer: Darreus Hsiao’s major is Dentistry. The first word of ‘Dentistry’ is ‘Dentistry’. The spelling of Dentistry is D, E, N, T, I, S, T, R, Y. The sequence of odd-positioned letters in ‘Dentistry’ is DNITY.
The answer is: DNITY
Table 1: Examples of the QA and reasoning tasks in Biography-Reasoning, associated with the Major type.
3.2 Controlled Study Design
To examine factual hallucinations caused by training with tasks containing new knowledge, we need to discriminate known and unknown knowledge, control their usage during training, and evaluate related hallucinations.
Since initially the model has no exposure to any knowledge of our synthetic dataset, we prepare the study by continue-pretraining the model with a subset of the knowledge, which becomes known to the model; and keep another subset of the knowledge as unknown. By mixing the constructed questions from known and unknown-knowledge in varying proportions, we are able to create situations where different proportion of newly introduced knowledge participates in training.
To evaluate how training leads to hallucinations, we reserve another subset of knowledge as test knowledge. The test knowledge are continue-pretrained together with the known knowledge during the preparation, but are kept away from further training. Therefore, the difference in performance on test set before and after training indicates the influence of factual hallucinations affected by training. In addition, we use the real-world ENTITYQUESTIONS dataset (Sciavolino et al. 2021) derived from Wikidata (Vrandečić and Krötzsch 2014) (denoted as wiki) as an out-of-distribution (OOD) test set to provide a more robust evaluation.
3.3 Models and Setups
We conduct experiments primarily using the Qwen2.5-1.5B model (Team 2024). As supplementary validation, we also perform partial experiments on Qwen3-8B (Team 2025) and Llama3.2-1B (Grattafiori et al. 2024) to assess generalization across model scales and architectures, with their results provided in Appendix G.
As our experiments are conducted on base models, we first apply SFT to endow them with the ability to answer questions in the evaluation sets. For QA analysis, SFT is conducted solely on knowledge QA data, whereas for reasoning, the model is trained jointly on both task types to ensure general reasoning competence.
All experiments are performed with full-parameter fine-tuning. Detailed hyperparameters are provided in the Appendix B. In the SFT phase, we default to training for 3 epochs, but we also provide results for training 1, 5, and 20 epochs in Appendix H. The settings of 1, 3, and 5 epochs simulate typical training schedules in practice, whereas 20 epochs allow the model to acquire most of the knowledge in the training set, even for previously unknown information.
Following Allen-Zhu and Li (2024) and Gekhman et al. (2024), we adopt Exact Match (EM) as the metric for both knowledge QA tasks and reasoning tasks to assess the accuracy of the final answers. Given that all test knowledge are known to the model, and the training and testing formats are consistent, there are no cases where the answers are correct but incorrectly formatted. We report the standard deviation of accuracy where applicable.
4 Hallucination Analysis
Using the Biography-Reasoning dataset, we conduct a systematic study on factual hallucinations induced by learning different tasks containing various types of new knowledge through SFT.
4.1 Knowledge QA
In this section, we analyze the impact of training on new knowledge in QA tasks. The baseline model is trained on samples constructed from the known knowledge of all four types. We then replace the knowledge of one entire type with unknown samples while keeping the other three types unchanged, resulting in four variant models. For each variant, we evaluate performance on three groups of QA test sets: (1) Same-Type QA (STQA): the test set corresponding to the replaced knowledge type; (2) Different-Type QA (DTQA): test sets of the remaining three types; (3) wiki: the real-world QA test set.
STQA DTQA wiki
-56.40 ± 4.28 -1.06 ± 0.30 -9.17 ± 4.17
Table 2: Average performance degradation (%, mean ± std) of four model variants, measured against the baseline model on different QA test groups. Detailed numerical results are reported in Appendix C.
Learning new knowledge induces factual hallucinations within the same type, with some spillover effects to other types. Table 2 presents the performance drop averaged across the four variant models. Training on unknown knowledge leads to substantial performance drops on the STQA test set, reducing the accuracy by more than half. We also observe cross-type degradation, as training on one type negatively impacts average performance on others, including the real-world wiki test set containing purely OOD knowledge. This confirms that learning new knowledge can induce hallucinations even on unrelated knowledge. Notably, the performance drop on DTQA is smaller than on wiki, as the former consists entirely of known data in the training set, which greatly mitigates the effect.
We further investigate how varying the proportion of unknown knowledge within a single type influences hallucination tendencies. Starting from the fully known-knowledge baseline, we progressively replace 5%, 10%, 20%, 50%, 80%, and 100% of the knowledge in one type with its corresponding unknown knowledge, while still keeping the other three types entirely known. Two strategies are considered for handling the remaining known knowledge within the modified type: KeepKnown, where the remaining known instances are retained, and RemoveKnown, where they are excluded from training.
Figure 2: Performance under two settings with different proportions of unknown knowledge in the same type and wiki test set.
As shown in Figure 2, the results across the four subplots are mutually corroborative, revealing a consistent pattern: the higher the proportion of unknown knowledge, the more severe the hallucination. In KeepKnown, performance on both the in-type test set and the OOD wiki test set degrades gradually at first, followed by a sharp decline as the amount of unknown knowledge increases. While RemoveKnown exhibits a more rapid degradation: within the target type, accuracy drops nearly linearly with the ratio of unknown samples; on the wiki set, performance declines sharply at low unknown counts and then plateaus. For the same replacement ratio, the key difference between KeepKnown and RemoveKnown lies in whether the type still contains known knowledge. We observe that this distinction has a substantial impact on model performance: RemoveKnown consistently underperforms KeepKnown. These observations suggest that sparse but fully unknown types are more disruptive than those containing a mixture of known and unknown knowledge, which differs from previous common understanding and poses new challenges for hallucination mitigation.
4.2 Knowledge-based Reasoning
For reasoning-related experiments, we train the model on both reasoning and QA tasks to facilitate a more reliable evaluation across both test sets. The baseline model is trained with all samples constructed from known knowledge. We then replace one reasoning task with instances derived from unknown knowledge and keep all other unchanged, resulting in 12 variant models.
We investigate how training on a knowledge-based reasoning task with unknown knowledge affects performance across different downstream tasks. Specifically, we examine six distinct test groups: (1) Same-Type Same-Reasoning (STSR): the exact reasoning task that trained with unknown knowledge type; (2) Same-Type Different-Reasoning (STDR): different reasoning tasks within the same knowledge type; (3) Different-Type Different-Reasoning (DTDR): all other reasoning tasks with different knowledge; (4) Same-Type QA (STQA): the QA task with the same knowledge type as STSR; (5) Different-Type QA (DTQA): QA tasks with other three knowledge types; and (6) wiki: the real-world QA test set.
We measure the relative performance change with respect to the baseline and compute the average difference within each of the six task groups. Results in Figure 3 show that learning reasoning tasks with new knowledge consistently induces performance degradation across all six groups. The overall trend aligns with previous findings: the most severe hallucinations occur in STSR, indicating strong intra-task interference, while hallucination on other task groups are relatively minor. A notable difference is that among other tasks, QA test sets exhibit even stronger hallucinations than seemingly more related reasoning tasks: STQA, DTQA, and even wiki show greater degradation than STDR and DTDR. Within QA, the degradation is more pronounced in STQA than in DTQA, implying that shared factual grounding heightens vulnerability.
Figure 3: The impact of learning new knowledge in reasoning tasks on the average performance across different groups. Results of each variant model on each dataset are presented in Appendix C.
5 KnownPatch: Mitigating Hallucination
In Section 4.1, we find that when all factual knowledge in a certain type is new, even if it constitutes a small proportion of the overall data, it significantly increases the risk of hallucination.
Prior work typically attempts to mitigate hallucinations by filtering out all data that contain unknown knowledge. However, in real-world settings, factual knowledge spans multiple categories and is intertwined with diverse task contexts, making it difficult to separate knowledge from task-related information. Identifying which portions of the data involve unknown knowledge is often ambiguous, and discarding entire samples that might contain it can be highly inefficient. As a result, perfect filtering is neither feasible nor effective in practice, and the remaining new knowledge can still induce serious hallucinations.
Even if perfect filtering were achievable, one might expect that transferring the new knowledge to Continued Pre-Training (CPT) would allow the model to first acquire it safely before SFT. However, our experiments show that learning new knowledge during the CPT phase can also cause hallucination behaviors similar to the observations in Section 4.1, with detailed results provided in Appendix E.
To address this issue, we propose KnownPatch, a simple yet effective approach that stabilizes the model by briefly reinforcing known knowledge at the final stage of training.
5.1 Intuition
We observe that learning tasks with known knowledge does not introduce instability, whereas training on unknown knowledge can easily disrupt the model’s behavior. Based on this observation, KnownPatch places a small portion of known knowledge at the very end of the training sequence. The idea is to allow the model, after being inevitably disturbed by new knowledge, to recover a stable state through a brief reinforcement of previously known information.
This approach is lightweight and easy to apply, as it only requires identifying a small subset of known samples from the SFT corpus. While precise separation of all known and unknown knowledge is practically infeasible, extracting a small subset of clearly known data is straightforward (Gekhman et al. 2024), making KnownPatch simple yet effective in stabilizing the overall training process.
5.2 Setups
KnownPatch modifies the training sequence by placing a small subset of SFT data with all known knowledge at the end of the training. In our setting, the data trained prior to this subset consists entirely of unknown knowledge across all of the types, to simulate the worst situation.
We consider two scenarios: (1) the ideal scenario, where the injected known knowledge covers all knowledge types that are involved during the previous training; (2) a more realistic scenario, where the injected known knowledge does not cover all types, and we specifically examine the case where known data from only one type is missing.
We evaluate KnownPatch across injection ratios of 5%, 10%, and 20% under both scenarios. The baseline model is trained on randomly shuffled data, where most knowledge within each type remains unknown; the theoretical upper bound is a model trained solely on known knowledge with the same dataset size.
To present the observations more clearly in the main text, we average and aggregate certain results, while detailed results are provided in Appendix F. In the first scenario, we average the results across the four test types (B, D, M, U) and collectively refer to them as QA, while the OOD wiki results, which are not involved in training, are reported separately. In the second scenario, we train four variant models, each corresponding to a different missing type in the injected known knowledge. We then average the performance of each variant on its respective missing-type test set and refer to this average as QA, while the averaged wiki performance across the four variants is reported separately.
5.3 Results
Figure 4: Averaged performance and attention score changes under different settings. The two bar groups on the left represent the performance of KnownPatch in the first scenario, while the bar group on the right corresponds to the second scenario. Values here represent the performance change percentage compared to the upper-bound model.
KnownPatch effectively mitigates factual hallucinations with minimal data injection. The left two groups of bars in Figure 4 show the performance of KnownPatch in knowledge QA tasks when injecting 5% and 20% of known data at the end of training. When the injected data covers all previous unknown knowledge types, KnownPatch consistently outperforms the baseline with randomly shuffled data across all injection ratios. Even with only 5% injected known data, KnownPatch recovers a substantial portion of the performance lost due to hallucinations. At 20%, the model’s performance becomes very close to that of the upper-bound model trained entirely on known knowledge. Notably, on the OOD wiki test set, performance with just 5% injected known data already approaches the upper bound, and at 20%, it even exceeds the performance of the fully known model, suggesting that late-stage injection of diverse known facts may enhance generalization and stabilize cross-domain knowledge application.
Even when one knowledge type is missing from the injection data, KnownPatch still substantially mitigates hallucination. As shown in the right group of bars in Figure 4, the model achieves significant performance gains even on the uncovered type, despite never being trained on its known facts. On the OOD wiki test set, performance also shows a clear improvement over the shuffled baseline and even comparable to the all-known upper-bound. This suggests that KnownPatch induces a global stabilization effect that benefits even types not directly represented in the patch.
KnownPatch also performs well in reasoning tasks. In knowledge-based reasoning tasks, when the injected patch data covers all previously unknown knowledge-based reasoning tasks, KnownPatch also achieves consistent performance improvements across all task types with different injection ratios. This effect can also transfer to QA tasks that are affected but not directly injected. Figure 5 shows the results of 20%, demonstrating the general effectiveness of the KnownPatch method.
Figure 5: Performance of KnownPatch on reasoning tasks when injecting 20% known data. The values here represent the accuracy percentage of this model compared to the fully known upper-bound model.
6 Interpretability with Attention Analysis
In this section, we analyze the mechanisms under new-knowledge-induced factual hallucinations through the attention scores. We measure the relative changes in entity attention and task performance in models trained with different data compositions (using the model trained with entirely known data as the baseline). We also analyze the attention pattern changes after applying the KnownPatch method.
6.1 Analysis Setup
We analyze the changes in attention scores that the model assigns to key entities in the context when applying knowledge. In the Biography-Reasoning dataset, the key entity in each question is the person name. Therefore, we examine the model’s attention score on the names when generating the first token of the knowledge.
Prior interpretability studies suggest that the model processes inputs in three stages: understanding, execution, and output, among which language-agnostic knowledge retrieval and abstract reasoning primarily occur in the middle-to-later layers (Wendler et al. 2024; Zhao et al. 2024). We also examine the model’s average attention to entities across different layers in both reasoning and knowledge QA tasks (detailed in Appendix D), with results shown in Figure 6, demonstrating that this attention is significantly higher in layers 12–24 (out of 28 total layers in Qwen2.5-1.5B), consistent with prior works. Therefore, in this study, we compute the average attention scores over layers 12 to 24 for interpretability analysis. We measure the relative change in entity attention to assess how training on new knowledge affects the model’s focus on critical entities, by comparing models trained under some new data to the model trained entirely on known knowledge, i.e. the upper bound model.
Figure 6: Attention score on the target name across layers in QA and reasoning training setups. The solid curves show the average attention score at each layer, aggregated across all datasets and instances. The shaded regions represent the standard deviation.
We use this method to analyze (1) the performance degradation of QA tasks under different proportions of unknown knowledge (Figure 2), (2) the impact of learning new knowledge in reasoning tasks (Figure 3), and (3) the effect of applying KnownPatch (Figure 4).
6.2 Analysis Results and Observations
Figures 7, 8 and 9 show the changes in entity attention and task performance for models trained on unknown knowledge in various settings.
Hallucinations emerge alongside a concurrent decline in entity attention and performance, and this coupling effect intensifies in the absence of known knowledge. Figure 7 presents the interpretability analysis corresponding to Figure 2. As the proportion of unknown instances within a knowledge type increases, the model’s attention to key entities gradually declines, accompanied by more severe hallucinations. The decline in entity attention under RemoveKnown is faster and leads to a more abrupt performance drop than under KeepKnown, indicating that the absence of known information within a knowledge type accelerates attention decay, which is strongly correlated with performance degradation.
Figure 7: Accuracy and attention score changes with different unknown data ratio in certain type.
Contextual similarity plays a crucial role in how hallucinations propagate across tasks. Figure 8 presents the interpretability analysis corresponding to Figure 3. We observe that both the performance drop and the reduction in attention to key entities are most pronounced in STSR, followed by the QA tasks. Overall, these two metrics exhibit a strong correlation across different task groups. Table 3 further shows the contextual similarity between STSR and other task groups. On average, QA tasks demonstrate the highest contextual similarity to the reasoning task involving unknown knowledge. Owing to their relatively long input contexts, reasoning tasks tend to be less similar to each other, whereas knowledge QA contexts often overlap with substrings of reasoning contexts. Taken together, these findings indicate that performance degradation grows with contextual similarity, implying that inter-task hallucinations are likely to spread through shared contextual patterns. Consequently, learning new knowledge in reasoning tasks is less likely to induce hallucinations in other reasoning tasks, while QA tasks are more vulnerable to such interference.
Figure 8: Accuracy and attention score changes compared to the all-known baseline model when learning new knowledge in reasoning tasks.
STSR STDR DTDR STQA DTQA wiki
1.0000 0.6164 0.5896 0.7312 0.6982 0.7199
Table 3: Averaged contextual similarity between STSR and other test groups, where each value denotes the mean similarity between the STSR context and all task contexts within the corresponding group. Similarity between A and B is computed as the proportion of tokens in B appearing in A.
KnownPatch stabilizes the model’s attention patterns and mitigates hallucinations. Figure 9 shows the performance drop and changes in attention on QA test sets relative to the all-known baseline model, providing an interpretability analysis corresponding to the right group of bars in Figure 4. It can be observed that KnownPatch can effectively restore the model’s focus on key entities in the questions, thereby significantly alleviating the hallucination effects caused by learning new knowledge.
Figure 9: Performance and attention score changes when learning new knowledge in QA tasks, and after applying KnownPatch (with 20% known data). QA represents the average across the four QA test sets, and error bars indicate standard deviations.
7 Conclusion
In this work, we present a systematic study on hallucinations caused by learning new knowledge in LLMs, examining their behavior across knowledge types and task types. Our experiments reveal that even a small number of fully unknown facts can trigger severe hallucinations, not only in the target task but also in other unrelated knowledge QA tasks, indicating a non-trivial inter-task impact. To address this hallucination, we propose KnownPatch, a simple but practical mitigation strategy that injects a small proportion of known knowledge samples during the final phase of fine-tuning. Without extensive data, KnownPatch effectively reduces hallucination across both QA and reasoning tasks even under partial type coverage. Through further analysis, we find that learning new knowledge shifts the model’s focus from key entities to other context, promoting erroneous binding and hallucination propagation. And this altered attention pattern can propagate along similar contexts. KnownPatch counteracts this by restoring entity-centric attention patterns.
Appendix A Dataset Details
A.1 Wiki Details
The ENTITYQUESTIONS dataset from Wikidata is divided into multiple subsets, such as P17, P20, etc., with each subset containing questions of the same format. For example, an instance from P17 is “Which country is Juniper Bank located in?”, and an instance from P20 is “Where did Connee Boswell die?”. Based on Gekhman et al. (2024)’s classification of knowledge, we categorize Wikipedia knowledge into four levels: HighlyKnown, MaybeKnown, WeaklyKnown, and Unknown. We construct the test set using subsets of HighlyKnown and MaybeKnown instances. To ensure the balanced distribution of the test set, we sample approximately the same number of questions from each subset, resulting in a final test set of 1,000 questions.
A.2 Biography-Reasoning Details
Each individual in the dataset is assigned four attributes: birth year, death year, major, and university. The dataset contains 3,000 individuals in total. Among them, 1,000 are kept as the unknown subset, while the remaining 2,000 individuals are trained during a CPT stage. Within the CPT subset, 1,000 individuals are reserved for building the test sets, and the other 1,000 are used as known knowledge to construct the training data. The detailed procedures for constructing names, attributes, and reasoning tasks are described below.
Names
The first name and last name of each individual are selected from separate pools and are ensured to be unique. For first names, we use 3,000 English names with an equal split between male and female names (this affects the use of gendered pronouns in reasoning tasks). For last names, we select 250 Chinese surnames , which are then randomly paired with the first names in a balanced manner. This random combination of English first names and Chinese last names is designed to generate synthetic individuals that minimize overlap with real-world knowledge already known to language models.
Attributes
The birth year of each synthetic individual is a random integer between 1800 and 1980. The death year is randomly assigned within the range of birth_year+30birth\_year+30 to min(2020,birth_year+100)min(2020,birth\_year+100), ensuring realistic lifespans. The major and university attributes are based on real-world entities. There are 50 universities in total, distributed across 10 countries (5 universities per country). There are also 50 majors, categorized into 10 broad fields (5 majors per field), e.g., Computer Science → Engineering.
CPT Data
The CPT data are mainly constructed in the form of biography texts. Here is an example of a biography:
Hannalee Sui was registered as born in 1974. Hannalee Sui brought her life to a close in 2015. Hannalee Sui participated in Accounting-related research. Hannalee Sui was officially registered at University of Alberta.
For the biographies used to construct known knowledge, each biography is rephrased 50 times to ensure consistent exposure. For those used to construct the test set, the biographies are divided into 10 subgroups, each rephrased 5, 10, …, up to 50 times, respectively. This design simulates a more realistic and diverse distribution of knowledge familiarity, reflecting varying degrees of knowledge internalization in practice.
Auxiliary Knowledge
To construct knowledge reasoning tasks, we introduce a set of auxiliary knowledge. Specifically, our dataset involves relations such as major → field (e.g., Computer Science belongs to Engineering) and university → country (e.g., Stanford University belongs to the United States). These auxiliary facts already exist in the model’s pre-trained knowledge base. To ensure that the model reliably retains them, we also rephrase each auxiliary fact 50 times and include them in the CPT data. All auxiliary knowledge is provided in Tables 5 and 6.
Reasoning Tasks
For each attribute of a synthetic individual, we construct three types of reasoning questions. In Table 4, we provide an example for each category of QA and reasoning questions in the dataset.
Each CR task involves two attributes: the primary attribute of interest and another randomly selected one. For major- and university-related CR tasks, which take a binary (Yes/No) form, we further constrain the sampling process to maintain an approximately balanced ratio of positive and negative instances (50% each).
During the SFT stage for reasoning tasks, we split the known individual set into two subsets: 80% are used to generate reasoning questions, and the remaining 20% are used for QA tasks. This allows the model to be exposed to both task types, ensuring a more reliable evaluation across reasoning and QA.
Category Example
B_QA Question: When was Darreus Hsiao born?
Answer: 1974
D_QA Question: When did Darreus Hsiao die?
Answer: 2017
M_QA Question: What major did Darreus Hsiao study?
Answer: Dentistry
U_QA Question: Which university did Darreus Hsiao graduate from?
Answer: Zhejiang University
B_SR Question: Is the number of Darreus Hsiao’s birth year an odd number?
Answer: Darreus Hsiao was born in 1974. 1974 % 2 = 0. So 1974 is not an odd number. The answer is: NO
B_CR Question: How many years apart is the birth year between Darreus Hsiao and Aydn Cheung?
Answer: Darreus Hsiao was born in 1974. Aydn Cheung was born in 1858. The difference is abs(1974 - 1858) = 116.
The answer is: 116
B_NR Question: What is the MScore of Darreus Hsiao’s birth year?
Answer: Darreus Hsiao was born in 1974. The four numbers are 1, 9, 7 and 4. So the MScore of it is 1 * 9 * 7 * 4 = 252.
The answer is: 252
D_SR Question: What year is the 10th anniversary of Darreus Hsiao’s death?
Answer: Darreus Hsiao died in 2017. 10 years after it should be 2017 + 10 = 2027.
The answer is: 2027
D_CR Question: Who died first, Darreus Hsiao or Aydn Cheung?
Answer: Darreus Hsiao died in 2017. Aydn Cheung died in 1919. 1919 is earlier than 2017. So Aydn Cheung died first.
The answer is: Aydn Cheung
D_NR Question: What is the AScore of Darreus Hsiao’s death year?
Answer: Darreus Hsiao died in 2017. The four numbers are 2, 0, 1 and 7. So the AScore of it is 2 + 0 + 1 + 7 = 10.
The answer is: 10
M_SR Question: What field does Darreus Hsiao’s major belong to?
Answer: Darreus Hsiao’s major is Dentistry. Dentistry belongs to Medicine.
The answer is: Medicine
M_CR Question: Do Darreus Hsiao and Virgus Hong’s majors belong to the same field?
Answer: Darreus Hsiao’s major is Dentistry. Dentistry belongs to Medicine. Virgus Hong’s major is Nursing. Nursing belongs to Medicine. Medicine and Medicine are the same.
The answer is: YES
M_NR Question: What is the sequence of odd-positioned letters in the first word of Darreus Hsiao’s major name?
Answer: Darreus Hsiao’s major is Dentistry. The first word of ‘Dentistry’ is ‘Dentistry’. The spelling of Dentistry is D, E, N, T, I, S, T, R, Y. The sequence of odd-positioned letters in ‘Dentistry’ is DNITY.
The answer is: DNITY
U_SR Question: In which country did Darreus Hsiao attend university?
Answer: Darreus Hsiao was graduated from Zhejiang University. Zhejiang University is located in China.
The answer is: China
U_CR Question: Are Darreus Hsiao and Angee Fung college alumni?
Answer: Darreus Hsiao was graduated from Zhejiang University. Saritha Tong was graduated from Kyoto University. Zhejiang University and Kyoto University are not the same.
The answer is: NO
U_NR Question: What is the sequence of the first and last letters of each word in Darreus Hsiao’s university name?
Answer: Darreus Hsiao was graduated from Zhejiang University, which can be splitted into words: Zhejiang, University. The first and last letters of ‘Zhejiang’ are ZG. The first and last letters of ‘University’ are UY. So, the whole sequence is ZGUY.
The answer is: ZGUY
Table 4: Examples of each QA and reasoning tasks in Biography-Reasoning.
Appendix B Training Details
In all CPT experiments, unless otherwise specified, we use a batch size of 16, a learning rate of 1e-5, a cutoff length of 512, and train for 1 epoch. In all SFT experiments (including knowledge QA and knowledge-based reasoning tasks), unless otherwise specified, we use a batch size of 32, a learning rate of 1e-5, and train for 3 epochs. We also did experiments of training 1 or 5 epochs, the results are presented in Appendix H. All experiments are conducted on up to 4 NVIDIA A6000 GPUs. The CPT stage is performed using LLaMA-Factory (Zheng et al. 2024).
Field Major
Economics Finance, Investment, Taxation, Insurance, Digital Economy
Law Intellectual Property, Criminal Justice, Sociology, International Politics, Diplomacy
Literature Journalism, Advertising, English, French, Russian
History Chinese History, World History, Museum Studies, Science History, Historical Geography
Science Mathematics, Physics, Chemistry, Biology, Geology
Engineering Computer Science, Software Engineering, Automation, Architecture, Electrical Engineering
Medicine Clinical Medicine, Dentistry, Pharmacy, Nursing, Public Health
Agriculture Agronomy, Horticulture, Plant Protection, Animal Science, Forestry
Management Accounting, Finance Management, Library Science, Tourism Management, Logistics Management
Art Fine Arts, Music, Dance, Art Theory, Environmental Design
Table 5: Auxiliary knowledge related to majors.
Country Universities
United States Harvard University, Stanford University, Princeton University,
Yale University, Columbia University
United Kingdom University of Oxford, University of Cambridge, Imperial College London,
University College London, University of Manchester
Canada University of Toronto, McGill University, University of Alberta,
McMaster University, University of Waterloo
Australia University of Melbourne, University of Sydney, University of Queensland,
Monash University, Macquarie University
Germany Heidelberg University, RWTH Aachen University, University of Freiburg,
University of Hamburg, University of Tübingen
France Sorbonne University, University of Paris, University of Strasbourg,
University of Lyon, University of Bordeaux
China Tsinghua University, Peking University, Fudan University,
Zhejiang University, Nanjing University
Japan Kyoto University, Osaka University, Tohoku University,
Nagoya University, Hokkaido University
Singapore Nanyang Technological University, Singapore Management University, Temasek Polytechnic,
Republic Polytechnic, Singapore Polytechnic
South Korea Seoul National University, Korea University, Yonsei University,
Sungkyunkwan University, Hanyang University
Table 6: Auxiliary knowledge related to universities.
Dataset Birth Death Major University
SR CR NR SR CR NR SR CR NR SR CR NR
$\text{All}_{\text{k}}$ 0.777 0.335 0.611 0.677 0.914 0.710 0.773 0.776 0.707 0.724 0.777 0.653
$\text{B\_SR}_{\text{unk}}$ 0.643 0.321 0.589 0.665 0.908 0.707 0.777 0.784 0.688 0.728 0.777 0.636
$\text{B\_CR}_{\text{unk}}$ 0.797 0.088 0.618 0.663 0.913 0.694 0.786 0.788 0.695 0.718 0.768 0.635
$\text{B\_NR}_{\text{unk}}$ 0.781 0.329 0.367 0.663 0.913 0.694 0.780 0.794 0.702 0.726 0.774 0.647
$\text{D\_SR}_{\text{unk}}$ 0.785 0.331 0.609 0.091 0.909 0.692 0.785 0.793 0.709 0.723 0.772 0.649
$\text{D\_CR}_{\text{unk}}$ 0.781 0.320 0.592 0.656 0.849 0.691 0.772 0.787 0.701 0.718 0.760 0.645
$\text{D\_NR}_{\text{unk}}$ 0.779 0.327 0.598 0.657 0.904 0.330 0.785 0.790 0.707 0.726 0.772 0.633
$\text{M\_SR}_{\text{unk}}$ 0.773 0.342 0.601 0.671 0.912 0.701 0.603 0.799 0.698 0.722 0.766 0.637
$\text{M\_CR}_{\text{unk}}$ 0.769 0.324 0.606 0.667 0.915 0.703 0.793 0.573 0.722 0.730 0.765 0.607
$\text{M\_NR}_{\text{unk}}$ 0.788 0.332 0.614 0.664 0.914 0.709 0.790 0.797 0.141 0.725 0.766 0.631
$\text{U\_SR}_{\text{unk}}$ 0.798 0.330 0.616 0.670 0.905 0.707 0.780 0.790 0.703 0.288 0.782 0.650
$\text{U\_CR}_{\text{unk}}$ 0.789 0.329 0.611 0.663 0.915 0.704 0.784 0.796 0.706 0.742 0.562 0.663
$\text{U\_NR}_{\text{unk}}$ 0.793 0.344 0.618 0.669 0.908 0.701 0.781 0.784 0.701 0.731 0.784 0.156
Table 7: Impact on other reasoning test sets when training new knowledge in reasoning tasks.
Dataset B_QA D_QA M_QA U_QA wiki
$\text{All}_{\text{k}}$ 0.578 0.665 0.297 0.673 0.286
$\text{B\_SR}_{\text{unk}}$ 0.562 0.651 0.316 0.668 0.289
$\text{B\_CR}_{\text{unk}}$ 0.581 0.639 0.328 0.684 0.275
$\text{B\_NR}_{\text{unk}}$ 0.568 0.616 0.165 0.671 0.279
$\text{D\_SR}_{\text{unk}}$ 0.569 0.669 0.279 0.670 0.274
$\text{D\_CR}_{\text{unk}}$ 0.552 0.627 0.293 0.670 0.273
$\text{D\_NR}_{\text{unk}}$ 0.563 0.658 0.318 0.681 0.283
$\text{M\_SR}_{\text{unk}}$ 0.573 0.663 0.319 0.669 0.282
$\text{M\_CR}_{\text{unk}}$ 0.566 0.603 0.157 0.598 0.272
$\text{M\_NR}_{\text{unk}}$ 0.577 0.545 0.190 0.655 0.266
$\text{U\_SR}_{\text{unk}}$ 0.578 0.571 0.210 0.606 0.290
$\text{U\_CR}_{\text{unk}}$ 0.564 0.657 0.355 0.685 0.276
$\text{U\_NR}_{\text{unk}}$ 0.565 0.619 0.299 0.683 0.284
Table 8: Impact on other QA test sets when training new knowledge in reasoning tasks.
Appendix C Detailed Results of Main Text
In this section, we detail the test results of each model on each dataset as shown in Table 2 and Figure 3 in the main text. In all settings, $\text{All}_{\text{k}}$ denotes the baseline model trained on data constructed entirely from known knowledge, while $\text{X}_{\text{unk}}$ refers to the variant where the subset X is replaced with unknown knowledge. Table 9 present the detailed results of the four variants of Table 2. Table 7 and Table 8 presents the twelve variants of reasoning tasks.
Model B_QA D_QA M_QA U_QA wiki
$\text{All}_{\text{k}}$ 0.549 0.609 0.546 0.464 0.199
$\text{B}_{\text{unk}}$ 0.200 0.591 0.555 0.466 0.195
$\text{D}_{\text{unk}}$ 0.539 0.225 0.531 0.464 0.188
$\text{M}_{\text{unk}}$ 0.545 0.596 0.255 0.456 0.183
$\text{U}_{\text{unk}}$ 0.531 0.606 0.546 0.252 0.157
Table 9: Hallucination induced by SFT on different unknown knowledge types.
Appendix D Attention Layer Selection
In Figure 6, the two lines represent the layer-wise entity attention patterns of two models across multiple datasets. The “QA” line corresponds to a model trained on all known QA questions (the baseline model in Figure 9) , with attention averaged over entity tokens in five QA test sets. The “Reasoning” line represents a model trained on a mixture of all 12 known reasoning tasks and QA questions (the baseline model in Figure 8) , with attention averaged over entity tokens in the reasoning test sets across all reasoning types.
Appendix E CPT Results
We investigate hallucination in models during the CPT phase. Using QA questions constructed from the Biography-Reasoning dataset, we construct CPT data concatenated via the EOS token. Note that this is the second CPT, because the first injection of known knowledge have undergone one CPT process, as described in Section 3. Building on the experimental setup described in the Appendix B, we conduct the following ablation studies: (1) the original experimental setting; (2) reducing the batch size from 16 to 1; (3) shortening the cutoff length from 512 to 32; and (4) increasing the total training data volume by a factor of 10.
The models are then evaluated with 5-shot QA format. We adapt the knowledge categorization method from Gekhman et al. (2024), with minor modifications. Specifically, we prompt the model with 5 different 5-shot samples. If the model answers correctly in at least one case, it is classified as Known; if all answers are incorrect, it is classified as Unknown. This is because the selection and order of few shots can significantly affect the model’s performance (Lu et al. 2022; Zhao et al. 2021), and we need to rule out this influence. The results are shown in Tables 10, 11, 12 and 13.
Among all the results, except for Table 10, there are quite serious hallucination phenomena. By varying different experimental settings, we rule out all interference factors and found that the number of steps for parameter updates may be the only variable that influences the degree of hallucination when LLM learns new knowledge. Specifically, due to the limited size of our dataset, the model is updated for only a small number of steps, resulting in relatively mild hallucinations as shown in Table 10. However, no matter whether we reduce the batch size, decrease the cutoff length, or increase the data volume, as long as the number of update steps increases, the hallucination phenomenon will become more serious.
Model B_QA D_QA M_QA U_QA
$\text{All}_{\text{k}}$ 0.655 0.715 0.429 0.430
$\text{B}_{\text{unk}}$ 0.560 0.706 0.326 0.416
$\text{D}_{\text{unk}}$ 0.621 0.684 0.346 0.421
$\text{M}_{\text{unk}}$ 0.652 0.712 0.368 0.422
$\text{U}_{\text{unk}}$ 0.646 0.713 0.366 0.426
Table 10: Accuracy on test sets of models trained on different unknown knowledge types during CPT with the original setting.
Model B_QA D_QA M_QA U_QA
$\text{All}_{\text{k}}$ 0.419 0.477 0.356 0.417
$\text{B}_{\text{unk}}$ 0.019 0.463 0.255 0.368
$\text{D}_{\text{unk}}$ 0.402 0.070 0.292 0.375
$\text{M}_{\text{unk}}$ 0.398 0.485 0.018 0.364
$\text{U}_{\text{unk}}$ 0.422 0.505 0.422 0.084
Table 11: Accuracy on test sets of models trained on different unknown knowledge types during CPT with setting (2): batch size reduced to 1.
Model B_QA D_QA M_QA U_QA
$\text{All}_{\text{k}}$ 0.477 0.522 0.520 0.439
$\text{B}_{\text{unk}}$ 0.081 0.538 0.523 0.464
$\text{D}_{\text{unk}}$ 0.464 0.114 0.562 0.416
$\text{M}_{\text{unk}}$ 0.453 0.539 0.027 0.407
$\text{U}_{\text{unk}}$ 0.460 0.565 0.408 0.162
Table 12: Accuracy on test sets of models trained on different unknown knowledge types during CPT with setting (3): cutoff length reduced to 32.
Model B_QA D_QA M_QA U_QA
$\text{All}_{\text{k}}$ 0.817 0.843 0.545 0.664
$\text{B}_{\text{unk}}$ 0.130 0.810 0.535 0.657
$\text{D}_{\text{unk}}$ 0.792 0.191 0.589 0.704
$\text{M}_{\text{unk}}$ 0.801 0.831 0.013 0.488
$\text{U}_{\text{unk}}$ 0.800 0.828 0.333 0.014
Table 13: Accuracy on test sets of models trained on different unknown knowledge types during CPT with setting (4) dataset increased by a factor of 10.
Appendix F Supplementary Results of KnownPatch
Figure 10 is the result of KnownPatch on QA tasks with different injection ratios, in addition to Figure 4 in the main text; Figures 11 and 12 are results of KnownPatch on reasoning tasks with injection ratios of 5% and 10%, in addition to Figure 5 (results of injection ratio 5%) in the main text; Figures 13, 14 and 15 are detailed results of KnownPatch when one knowledge type is missed, with injection ratios 20%, 10% and 5%, respectively.
Figure 10: Performance of KnownPatch on QA task when injecting 5% (upper), 10% (middle) and 20% (lower) known data. All experiments trained for 3 epoch.
Figure 11: KnownPatch on reasoning tasks with 5% injection ratio. All experiments trained for 3 epoch.
Figure 12: KnownPatch on reasoning tasks with 10% injection ratio. All experiments trained for 3 epoch.
Figure 13: KnownPatch (missing one knowledge type) on QA tasks with an injection ratio of 20%. All experiments trained for 3 epoch.
Figure 14: KnownPatch (missing one knowledge type) on QA tasks with an injection ratio of 10%. All experiments trained for 3 epoch.
Figure 15: KnownPatch (missing one knowledge type) on QA tasks with an injection ratio of 5%. All experiments trained for 3 epoch.
Appendix G Results on Different Models
We used Qwen2.5-1.5B (main text) and Qwen3-8B, Llama3.2-1B (appendix), spanning architectures and sizes, all supporting our conclusions. Due to resource limitations, larger-scale training was infeasible. For larger models, prior work (Allen-Zhu and Li 2025) shows they retain factual knowledge better.
G.1 Results on Llama-3.2-1B
In this section we provide results of Llama-3.2-1B. Table 14 (similar to Table 2) provides the hallucination results in QA tasks when learning new knowledge; Figure 16 (similar to Figure 3) shows the impact of new knowledge in reasoning tasks on different groups.
Model B_QA D_QA M_QA U_QA wiki
$\text{All}_{\text{k}}$ 0.863 0.855 0.766 0.668 0.155
$\text{B}_{\text{unk}}$ 0.332 0.837 0.761 0.657 0.116
$\text{D}_{\text{unk}}$ 0.836 0.359 0.767 0.664 0.145
$\text{M}_{\text{unk}}$ 0.845 0.848 0.478 0.664 0.152
$\text{U}_{\text{unk}}$ 0.855 0.860 0.761 0.387 0.141
Table 14: Llama-3.2-1B model’s hallucination induced by training on different unknown knowledge types in QA tasks.
Figure 16: The impact of learning new knowledge in reasoning tasks on the average performance across different groups (on the Llama-3.2-1B model).
We also perform the same interpretability analysis as Section 6 and Appendix D on the Llama-3.2-1B model. Based on the results of Figure 17 (similar to Figure 6), we chose its layers 4-14 for further interpretability analysis, and Figure 18 (similar to Figure 9) shows the results.
Figure 17: Llama-3.2-1B model’s attention score on the target name across layers in QA and reasoning training setups. The solid curves show the average attention score at each layer, aggregated across all datasets and instances. The shaded regions represent the standard deviation.
Figure 18: Llama-3.2-1B model’s performance and attention score changes when learning new knowledge in QA tasks, and after applying KnownPatch (with 20% known data). QA represents the average across the four QA test sets, and error bars indicate standard deviations.
G.2 Results on Qwen3-8B
Due to the large scale of model parameters, our training setting differ from the default one. We only fine-tune for 1 epoch with a learning rate 5e-6 in all the SFT experiments.
Table 15 (similar to Table 2) provides the hallucination results in QA tasks when learning new knowledge; Figure 19 (similar to Figure 3) shows the impact of new knowledge in reasoning tasks on different groups.
Model B_QA D_QA M_QA U_QA wiki
$\text{All}_{\text{k}}$ 0.870 0.864 0.867 0.709 0.301
$\text{B}_{\text{unk}}$ 0.157 0.850 0.855 0.709 0.253
$\text{D}_{\text{unk}}$ 0.846 0.165 0.847 0.676 0.277
$\text{M}_{\text{unk}}$ 0.846 0.855 0.167 0.680 0.285
$\text{U}_{\text{unk}}$ 0.833 0.852 0.852 0.149 0.229
Table 15: Qwen3-8B model’s hallucination induced by training on different unknown knowledge types in QA tasks.
Figure 19: The impact of learning new knowledge in reasoning tasks on the average performance of different groups (on the Qwen3-8B model).
We perform the same analysis as Section 6 and Appendix D on the Qwen3-8B model. Based on the results of Figure 20 (similar to Figure 6), we chose its layers 9-27 for further interpretability analysis, and Figure 21 (similar to Figure 9) shows the results.
Figure 20: Qwen3-8B model’s attention score on the target name across layers in QA and reasoning training setups. The solid curves show the average attention score at each layer, aggregated across all datasets and instances. The shaded regions represent the standard deviation.
Figure 21: Qwen3-8B model’s performance and attention score changes when learning new knowledge in QA tasks, and after applying KnownPatch (with 20% known data). QA represents the average across the four QA test sets, and error bars indicate standard deviations.
Appendix H Results on Different Training Epochs
All results presented in the main text are obtained after SFT for 3 epochs. In this section, we report the results of the Qwen2.5-1.5B model under the same experimental configurations, with the number of training epochs adjusted to 1, 5 and 20. Notably, after 3 epochs of training, the model already achieves over 95% accuracy on questions derived from known knowledge in the training set, and about 50% accuracy on those constructed from unknown knowledge. When training is extended to 20 epochs, the model reaches over 95% accuracy on unknown knowledge questions in the training set, and further training brings little additional improvement. We observe that the overall trends and results remain consistent across different numbers of training epochs.
H.1 1 Epoch
Table 16 (similar to Table 2) provides the hallucination results in QA tasks when learning new knowledge; Figure 22 (similar to Figure 2) shows the performance after learning different proportions of unknown knowledge; Figure 23 (similar to Figure 3) shows the impact of new knowledge in reasoning tasks on different groups; Figure 24 (similar to Figure 10) reports the performance of KnownPatch on QA tasks when injecting 20% known data; Figure 25 (similar to Figure 5) reports performance of KnownPatch on reasoning tasks when injecting 20% known data; Figure 26 (similar to Figure 13) reports performance of KnownPatch when one knowledge type is missing with 20% injection ratio; Figure 27 (similar to Figure 8) reports the accuracy and attention score changes when learning new knowledge in reasoning tasks; Figure 28 (similar to Figure 7) reports the accuracy and attention score changes after learning different proportions of unknown knowledge; Figure 29 (similar to Figure 9) reports the performance and attention score changes before and after applying KnownPatch.
Model B_QA D_QA M_QA U_QA wiki
$\text{All}_{\text{k}}$ 0.548 0.568 0.475 0.458 0.198
$\text{B}_{\text{unk}}$ 0.198 0.527 0.481 0.448 0.199
$\text{D}_{\text{unk}}$ 0.518 0.280 0.471 0.442 0.190
$\text{M}_{\text{unk}}$ 0.517 0.576 0.193 0.423 0.173
$\text{U}_{\text{unk}}$ 0.529 0.594 0.467 0.304 0.174
Table 16: Hallucination induced by training on different unknown knowledge types in QA tasks. All experiments trained for 1 epoch.
Figure 22: Performance in QA tasks under two settings with different proportions of unknown knowledge in the same type and wiki test set. All experiments trained for 1 epoch.
Figure 23: The impact of learning new knowledge in reasoning tasks on the average performance of different groups. All experiments trained for 1 epoch.
Figure 24: Performance of KnownPatch on QA task when injecting 20% known data. All experiments trained for 1 epoch.
Figure 25: Performance of KnownPatch on reasoning task when injecting 20% known data. The value here represents the accuracy percentage of this model compared to the fully known baseline model. All experiments trained for 1 epoch.
Figure 26: KnownPatch (missing one knowledge type) on QA tasks with an injection ratio of 20%. All experiments trained for 1 epoch.
Figure 27: Accuracy and attention score changes when learning new knowledge in reasoning tasks. All experiments trained for 1 epoch.
Figure 28: Accuracy and attention score changes with different unknown data ratio in certain type in QA tasks. All experiments trained for 1 epoch.
Figure 29: Performance and attention score changes when learning new knowledge in QA tasks, and after applying KnownPatch (with 20% known data). QA represents the average across the four QA test sets, and error bars indicate standard deviations. All experiments trained for 1 epoch.
H.2 5 Epochs
Table 17 (similar to Table 2) provides the hallucination results in QA tasks when learning new knowledge; Figure 30 (similar to Figure 2) shows the performance after learning different proportions of unknown knowledge; Figure 31 (similar to Figure 3) shows the impact of new knowledge in reasoning tasks on different groups; Figure 32 (similar to Figure 10) reports the performance of KnownPatch on QA task with 20% injection ratio; Figure 33 (similar to Figure 5) reports performance of KnownPatch when injecting 20% known data; Figure 34 (similar to Figure 13) reports performance of KnownPatch when one knowledge type is missing when injecting 20% known data; Figure 35 (similar to Figure 8) reports the accuracy and attention score changes when learning new knowledge in reasoning tasks; Figure 36 (similar to Figure 7) reports the accuracy and attention score changes after learning different proportions of unknown knowledge; Figure 37 (similar to Figure 9) reports the performance and attention score changes before and after applying KnownPatch.
Model B_QA D_QA M_QA U_QA wiki
$\text{All}_{\text{k}}$ 0.529 0.600 0.552 0.447 0.200
$\text{B}_{\text{unk}}$ 0.234 0.573 0.550 0.434 0.194
$\text{D}_{\text{unk}}$ 0.505 0.289 0.533 0.435 0.181
$\text{M}_{\text{unk}}$ 0.523 0.606 0.215 0.432 0.171
$\text{U}_{\text{unk}}$ 0.523 0.607 0.554 0.245 0.144
Table 17: Hallucination induced by training on different unknown knowledge types in QA tasks. All experiments trained for 5 epoch.
Figure 30: Performance in QA tasks under two settings with different proportions of unknown knowledge in the same type and wiki test set. All experiments trained for 5 epoch.
Figure 31: The impact of learning new knowledge in reasoning tasks on the average performance of different groups. All experiments trained for 5 epoch.
Figure 32: Performance of KnownPatch on QA task when injecting 5% (upper), 10% (middle) and 20% (lower) known data. All experiments trained for 5 epoch.
Figure 33: Performance of KnownPatch on reasoning task when injecting 20% known data. The value here represents the accuracy percentage of this model compared to the fully known baseline model. All experiments trained for 5 epoch.
Figure 34: KnownPatch (missing one knowledge type) on QA tasks with an injection ratio of 20%. All experiments trained for 5 epoch.
Figure 35: Accuracy and attention score changes when learning new knowledge in reasoning tasks. All experiments trained for 5 epoch.
Figure 36: Accuracy and attention score changes with different unknown data ratio in certain type in QA tasks. All experiments trained for 5 epoch.
Figure 37: Performance and attention score changes when learning new knowledge in QA tasks, and after applying KnownPatch (with 20% known data). QA represents the average across the four QA test sets, and error bars indicate standard deviations. All experiments trained for 5 epoch.
H.3 20 Epochs
Table 18 (similar to Table 2) provides the hallucination results in QA tasks when learning new knowledge; Figure 38 (similar to Figure 2) shows the performance after learning different proportions of unknown knowledge; Figure 39 (similar to Figure 3) shows the impact of new knowledge in reasoning tasks on different groups; Figure 40 (similar to Figure 10) reports the performance of KnownPatch on QA task when injecting 20% known data; Figure 41 (similar to Figure 5) reports performance of KnownPatch when injecting 20% known data; Figure 42 (similar to Figure 13) reports performance of KnownPatch when one knowledge type is missing when injecting 20% known data; Figure 43 (similar to Figure 8) reports the accuracy and attention score changes when learning new knowledge in reasoning tasks; Figure 44 (similar to Figure 7) reports the accuracy and attention score changes after learning different proportions of unknown knowledge; Figure 45 (similar to Figure 9) reports the performance and attention score changes before and after applying KnownPatch.
Model B_QA D_QA M_QA U_QA wiki
$\text{All}_{\text{k}}$ 0.498 0.611 0.540 0.371 0.186
$\text{B}_{\text{unk}}$ 0.183 0.503 0.516 0.380 0.176
$\text{D}_{\text{unk}}$ 0.443 0.231 0.499 0.352 0.159
$\text{M}_{\text{unk}}$ 0.469 0.544 0.343 0.360 0.178
$\text{U}_{\text{unk}}$ 0.452 0.549 0.485 0.252 0.171
Table 18: Hallucination induced by training on different unknown knowledge types in QA tasks. All experiments trained for 20 epoch.
Figure 38: Performance in QA tasks under two settings with different proportions of unknown knowledge in the same type and wiki test set. All experiments trained for 20 epoch.
Figure 39: The impact of learning new knowledge in reasoning tasks on the average performance of different groups. All experiments trained for 20 epoch.
Figure 40: Performance of KnownPatch on QA task when injecting 5% (upper), 10% (middle) and 20% (lower) known data. All experiments trained for 20 epoch.
Figure 41: Performance of KnownPatch on reasoning task when injecting 20% known data. The value here represents the accuracy percentage of this model compared to the fully known baseline model. All experiments trained for 20 epoch.
Figure 42: KnownPatch (missing one knowledge type) on QA tasks with an injection ratio of 20%. All experiments trained for 20 epoch.
Figure 43: Accuracy and attention score changes when learning new knowledge in reasoning tasks. All experiments trained for 20 epoch.
Figure 44: Accuracy and attention score changes with different unknown data ratio in certain type in QA tasks. All experiments trained for 20 epoch.
Figure 45: Performance and attention score changes when learning new knowledge in QA tasks, and after applying KnownPatch (with 20% known data). QA represents the average across the four QA test sets, and error bars indicate standard deviations. All experiments trained for 20 epoch.

### v2

@@
-Understanding New-Knowledge-Induced Factual Hallucinations in LLMs: Analysis, Solution and Interpretation
+Understanding New-Knowledge-Induced Factual Hallucinations in LLMs: Analysis and Interpretation
 Abstract
-Previous studies show that introducing new knowledge during large language models (LLMs) fine-tuning can lead to the generation of erroneous output when tested on known information, thereby triggering factual hallucinations.
-However, existing studies have not deeply investigated the specific manifestations and underlying mechanisms of these hallucinations.
+Prior works have shown that fine-tuning on new knowledge can induce factual hallucinations in large language models (LLMs), leading to incorrect outputs when evaluated on previously known information.
+However, the specific manifestations of such hallucination and its underlying mechanisms remain insufficiently understood.
 Our work addresses this gap by designing a controlled dataset Biography-Reasoning, and conducting a fine-grained analysis across multiple knowledge types and two task types, including knowledge question answering (QA) and knowledge reasoning tasks.
-We find that when fine-tuned on a dataset in which a specific knowledge type consists entirely of new knowledge, LLMs exhibit significantly increased hallucination tendencies.
-This suggests that the high unfamiliarity of a particular knowledge type, rather than the overall proportion of new knowledge, is a stronger driver of hallucinations, and these tendencies can even affect other knowledge types in QA tasks.
-To mitigate such factual hallucinations, we propose KnownPatch, which patches a small number of known knowledge samples in the later stages of training, effectively alleviating new-knowledge-induced hallucinations.
-Through attention analysis, we find that learning new knowledge reduces the model’s attention to key entities in the question, thus causing excessive focus on the surrounding context, which may increase the risk of hallucination.
-Moreover, the attention pattern can propagate to similar contexts, facilitating the spread of hallucinations to textually similar questions.
-Our method effectively mitigates the disruption of new knowledge learning to the model’s attention on key entities, accompanied by improved performance.
-Codehttps://github.com/NJUNLP/New-Knowledge-Induced-Factual-Hallucinations
+We find that hallucinations not only severely affect tasks involving newly introduced knowledge, but also propagate to other evaluation tasks.
+Moreover, when fine-tuning on a dataset in which a specific knowledge type consists entirely of new knowledge, LLMs exhibit elevated hallucination tendencies.
+This suggests that the degree of unfamiliarity within a particular knowledge type, rather than the overall proportion of new knowledge, is a stronger driver of hallucinations.
+Through interpretability analysis, we show that learning new knowledge weakens the model’s attention to key entities in the input question, leading to an over-reliance on surrounding context and a higher risk of hallucination.
+Conversely, reintroducing a small amount of known knowledge during the later stages of training restores attention to key entities and substantially mitigates hallucination behavior.
+Finally, we demonstrate that disrupted attention patterns can propagate across lexically similar contexts, facilitating the spread of hallucinations beyond the original task.
+Understanding New-Knowledge-Induced Factual Hallucinations in LLMs:
+Analysis and Interpretation
+Renfei Dang1, Peng Hu1, Zhejian Lai1, Changjiang Gao1, Min Zhang2, Shujian Huang1 1 National Key Laboratory for Novel Software Technology, Nanjing University, Nanjing, China 2 Huawei Translation Services Center, Beijing, China {dangrf,hup,laizj,gaocj}@smail.nju.edu.cn, huangsj@nju.edu.cn, zhangmin186@huawei.com
+11footnotetext: Equal contribution.22footnotetext: Corresponding author.
 1 Introduction
-LLMs embed rich factual knowledge in their parameters during pre-training on massive text corpora (Petroni et al.
-2019; Cohen et al.
-2023).
-Subsequently, post-training enables them to learn to follow human instructions and exhibit superior performance across various downstream tasks (Ouyang et al.
-2022; Wei et al.
-2022).
+Large language models (LLMs) acquire rich factual knowledge during pre-training on massive text corpora (Petroni et al., 2019; Cohen et al., 2023), and are subsequently post-trained to follow human instructions and perform a wide range of downstream tasks (Ouyang et al., 2022; Wei et al., 2022).
 Figure 1: The impact of learning new knowledge on attention patterns and hallucination behavior.
-When an model is trained on unknown facts, it may be more prone to produce factual hallucinations, the severity of which correlates with the degree of attention paid to key entities and is modulated by contextual similarity.
-By injecting a small amount of known knowledge at the end of training via KnowPatch, this issue can be effectively mitigated.
+Training a model on new knowledge can induce factual hallucinations, whose severity is correlated with the model’s attention scores on key entities.
+Moreover, hallucinations are more likely to occur on test instances whose input contexts are lexically similar to the contexts of training tasks containing unknown knowledge.
+By injecting a small amount of known knowledge at the end of training, this issue can be effectively mitigated.
 However, during the Supervised Fine-Tuning (SFT) phase, models may encounter new knowledge not covered in pre-training.
-Prior researches (Ghosal, Hashimoto, and Raghunathan 2024; Lin et al.
-2023; Ovadia et al.
-2023; Gekhman et al.
-2024; Sun et al.
-2025) suggest that introducing new knowledge in the post-training phase increases the risk of factual hallucinations, where models generate fabricated yet plausible statements.
-This occurs because, when models learn new facts absent from pre-training, they may erroneously generate related information in irrelevant contexts (Gekhman et al.
-2024; Sun et al.
-2025).
-These studies primarily focus on the effects within knowledge-intensive QA tasks during SFT, and we advance this line of research by conducting a systematic analysis of fine-grained manifestations and underlying causes of hallucinations.
-To enable a comprehensive investigation, we construct a controlled experimental dataset, Biography-Reasoning.
-The dataset is composed of biographical entities and their four attributes (birth, death, major, and university), which serve as four knowledge types.
+Prior researches (Ghosal et al., 2024; Lin et al., 2023; Ovadia et al., 2023; Gekhman et al., 2024; Sun et al., 2025) suggest that introducing new knowledge in the post-training phase increases the risk of factual hallucinations, where models generate fabricated yet plausible statements.
+This occurs because, when models learn new knowledge, they may erroneously generate related information in irrelevant contexts (Gekhman et al., 2024; Sun et al., 2025).
+These studies primarily focus on the effects within knowledge-intensive QA tasks during SFT, and we advance this line of research by investigating the fine-grained manifestations and underlying causes of hallucinations.
+To support this investigation, we construct a controlled experimental dataset Biography-Reasoning.
+The dataset is composed of biographical entities and their four attributes, which serve as four knowledge types.
 We further design twelve reasoning tasks using these knowledge.
-By controlling the proportion of known and unknown knowledge within different types and tasks in the training data, we systematically investigate the impact of learning new knowledge on hallucination risks.
-Our experiments reveal that, for knowledge QA tasks, training on unknown knowledge significantly elevates hallucination risks in the same type test set, with some cross-type influence on other QA test sets; in tasks involving knowledge reasoning, training on reasoning tasks containing unknown knowledge primarily affects the same reasoning task, with some impact on knowledge QA and limited influence on other reasoning tasks.
-Importantly, we find that if a knowledge type consists entirely of new knowledge, even a small amount of such data can markedly increase hallucination tendencies.
-This poses challenges for non-perfect filtering-based hallucination mitigation methods: inadvertently retained small amounts of unknown data can trigger severe hallucinations.
-To address this challenge, we propose KnownPatch, a simple yet effective strategy that places a small number of known knowledge in later stages of training.
-Instead of relying on exhaustive filtering of unknown content, our method stabilizes learning by reinducing known knowledge late in training.
-Experiments show that even a small injection of known data significantly reduces hallucination tendencies across diverse settings.
-Finally, we perform an interpretability analysis of the model’s attention distribution.
-The results reveal that learning new knowledge significantly weakens the model’s attention to key entities in the question, instead may erroneously bind new knowledge to other contexts, thereby triggering factual hallucinations.
-Therefore, tasks that share similar contexts are more likely to be affected and exhibit stronger hallucination tendencies.
-KnownPatch effectively restores and enhances the model’s attention to key entities, thereby mitigating hallucination risks.
+By controlling the proportion of known and unknown knowledge within different types and tasks in the training data, we systematically analyze the impact of learning new knowledge on hallucination risks.
+Our experiments reveal that training on unknown knowledge significantly elevates hallucination risks in the same task, while also inducing non-negligible hallucination effects on other out-of-domain test tasks.
+Importantly, we further find that when a knowledge type consists entirely of new knowledge, even a small amount of such data can markedly increase hallucination tendencies.
+Through further interpretability analyses, we find that learning new knowledge significantly weakens the model’s attention to key entities in the question, thereby triggering factual hallucinations.
+In contrast, training on known knowledge strengthens the model’s attention to key entities.
+Motivated by this observation, we introduce a simple training method KnownPatch, which restores disrupted attention patterns by injecting a small amount of known knowledge during the later stages of training, and thus alleviates factual hallucinations.
+Finally, by constructing carefully designed variants of reasoning tasks, we demonstrate that lexical similarity (measured by token overlap between contexts), rather than semantic similarity of the contexts, is the primary driver of hallucination propagation across tasks.
 These findings are visually presented in Figure 1.
-In summary, the main contributions of this paper are:
+The main contributions of this paper are:
   • Fine-Grained Analysis: A detailed analysis across knowledge types and task types reveals the manifestations of new-knowledge-induced hallucinations, showing that when all knowledge within a specific type is entirely unknown, it is more likely to trigger severe hallucinations, even on unrelated QA test sets.
-  • Mitigation Strategy: The introduction of KnownPatch, a training method that places a small amount of known knowledge in the later stages, significantly reducing hallucination risks without data modifications.
-  • Mechanism Interpretability: An analysis of attention mechanisms shows that learning new knowledge reduces attention to key question entities and erroneously binds new knowledge to contexts, causing hallucinations.
-Moreover, similar contexts facilitate the spread of these altered attention patterns, enabling cross-task hallucination effects.
+  • Mechanism Interpretability: An analysis of attention mechanisms shows that learning new knowledge reduces attention to key question entities, causing hallucinations.
+In addition, lexically similar contexts facilitate the spread of these attention patterns, enabling cross-task hallucination effects.
 2 Related Work
 New Knowledge and Hallucinations
-Existing studies have indicated that introducing new knowledge into LLMs may trigger hallucinations (Ghosal, Hashimoto, and Raghunathan 2024; Lin et al.
-2023; Ovadia et al.
-2023).
-Subsequent works have provided deeper analyses of this phenomenon(Gekhman et al.
-2024; Kang et al.
-2024; Sun et al.
-2025).
+Existing studies have indicated that introducing new knowledge into LLMs may trigger hallucinations (Ghosal et al., 2024; Lin et al., 2023; Ovadia et al., 2023).
+Subsequent works have provided deeper analyses of this phenomenon (Gekhman et al., 2024; Kang et al., 2024; Sun et al., 2025).
 Gekhman et al.
@@
 (2024) analyzed that when fine-tuned LLMs encounter unknown queries during testing, their responses imitate those associated with unknown examples in the fine-tuning data.
-Sun et al.
-(2025), from the perspective of token probabilities, examined how, after learning new knowledge, the generation probability of relevant answer entity tokens in irrelevant contexts increases significantly, suggesting that the model may erroneously generalize knowledge, leading to hallucinations.
+From the perspective of token probabilities, Sun et al.
+(2025) show that after learning new knowledge, the generation probabilities of answer entity tokens increase significantly even in irrelevant contexts, suggesting that the model may over-generalize newly acquired knowledge and consequently produce hallucinations.
 However, previous studies focus mainly on closed-book QA settings with mixed knowledge types during training, while our controlled setup disentangles them to provide a more detailed analysis of new knowledge-induced hallucinations across types and tasks.
@@
 Numerous studies are currently exploring ways to mitigate model hallucinations.
-A common approach involves providing additional relevant context to the model to reduce hallucinations during generation, such as through retrieval from knowledge bases or leveraging other large models to generate context(Shuster et al.
-2021; Sun et al.
-2022; Asai et al.
-2024; Feng et al.
-2023).
-Additionally, some research explicitly avoids hallucination risks by refusing to answer uncertain or unfamiliar questions(Yadkori et al.
-2024; Zhu et al.
-2025; Duwal 2025).
-In another direction, many studies encourage the model to generate more known knowledge from pre-training, for example, by promoting factual outputs via reinforcement learning (Rafailov et al.
-2023; Kang et al.
-2024; Li and Ng 2025; Gu et al.
-2025) or by training only on known knowledge during supervised fine-tuning (Lin et al.
-2024; Ghosal, Hashimoto, and Raghunathan 2024; Liu et al.
-2024) to enhance the model.
+A common approach involves providing additional relevant context to the model to reduce hallucinations during generation, such as through retrieval from knowledge bases or leveraging other large models to generate context (Shuster et al., 2021; Sun et al., 2022; Asai et al., 2024; Feng et al., 2023).
+Additionally, some research explicitly avoids hallucination risks by refusing to answer uncertain or unfamiliar questions (Yadkori et al., 2024; Zhu et al., 2025; Duwal, 2025).
+In another direction, many studies encourage the model to generate more known knowledge from pre-training, for example, by promoting factual outputs via reinforcement learning (Rafailov et al., 2023; Kang et al., 2024; Li and Ng, 2025; Gu et al., 2025) or by training only on known knowledge during supervised fine-tuning (Lin et al., 2024; Ghosal et al., 2024; Liu et al., 2024) to enhance the model.
 Our work builds on SFT with known knowledge approach, but rather than pursuing comprehensive filtering across all training data, KnownPatch only introduces a small number of known knowledge samples in the later stages of training, and alleviates the model’s tendency for hallucination.
@@
 However, in real-world datasets, most factual knowledge may have already been seen by LLMs during pre-training, making it difficult to precisely control whether the knowledge being learned is new to the model.
-To address this limitation, we construct a controlled experimental environment and a synthetic dataset named Biography-Reasoning, which allows a controllable examination of hallucination behaviors under varying knowledge types and task types.
+To address this limitation, we construct a synthetic dataset named Biography-Reasoning, which allows a controllable examination of hallucination behaviors under varying knowledge types and task types.
 3.1 Biography-Reasoning Dataset
@@
 (2025), we design the Biography-Reasoning dataset.
-The dataset centers on individuals as the key entities, with each person associated with four attributes: birth year (B), death year (D), major (M), and university (U).
+The dataset centers on individuals as the key entities, with each person associated with four attributes: birth year, death year, major, and university.
 We refer to the same attribute of different individuals as a knowledge type.
-Our dataset includes two main types of knowledge-related tasks, i.e.
-knowledge QA and knowledge-based reasoning tasks.
+Our dataset includes two types of knowledge-related tasks: knowledge QA and knowledge-based reasoning tasks.
 For knowledge QA tasks, questions are formulated by directly querying one of the attributes given the person’s name.
-Each task consists of questions on a single type, resulting in four QA tasks (B_QA, D_QA, M_QA and U_QA).
-For knowledge-based reasoning tasks, we design three fundamental types of chain-of-thought-requiring reasoning tasks. Specifically, these include:
-  • SR (Single Reasoning): extracting one attribute from a single entity and performing a simple reasoning process;
-  • CR (Comparative Reasoning): extracting one attribute from each of two entities and performing comparative reasoning between them;
-  • NR (Novel Reasoning): extracting one attribute from a single entity and performing a newly defined reasoning task, such as mathematical or symbolic reasoning.
+Each task consists of questions on a single type, resulting in four QA tasks (e.g., Major_QA).
+For knowledge-based reasoning tasks, we design three types of chain-of-thought-requiring reasoning tasks. Specifically, these include:
+  • Single Reasoning: extracting one attribute from a single entity and performing a simple reasoning process;
+  • Comparative Reasoning : extracting one attribute from each of two entities and performing comparative reasoning between them;
+  • Novel Reasoning: extracting one attribute from a single entity and performing a newly defined reasoning task, such as mathematical or symbolic reasoning.
 Table 1 presents examples of the constructed questions.
-Some of the reasoning tasks are intentionally designed to be more complex than mere knowledge extraction as QA problems.
-They require further reasoning, as well as auxiliary knowledge (e.g., the Major Dentistry belongs to the field Medicine), which the model is expected to contain.
+The reasoning tasks are intentionally designed to be more complex than mere knowledge extraction as QA problems.
+Some of them require auxiliary knowledge (e.g., the major Dentistry belongs to the field Medicine), which the model is expected to contain.
 To further guarantee the model’s proficiency, we additionally collect and train on these auxiliary facts.
-For each knowledge type we construct one QA and three reasoning tasks, leading to a total of four QA and 12 reasoning tasks per individual. Further dataset details can be found in Appendix A.
+For each knowledge type we construct one QA and three reasoning tasks, leading to a total of 4 QA and 12 reasoning tasks per individual. Further dataset details can be found in Appendix A.
 Category Example
-M_QA Question: What major did Darreus Hsiao study?
+QA Question: What major did Darreus Hsiao study?
 Answer: Dentistry
-M_SR Question: What field does Darreus Hsiao’s major belong to?
+Single Reasoning Question: What field does Darreus Hsiao’s major belong to?
 Answer: Darreus Hsiao’s major is Dentistry. Dentistry belongs to Medicine.
 The answer is: Medicine
-M_CR Question: Do Darreus Hsiao and Virgus Hong’s majors belong to the same field?
+Comparative Reasoning Question: Do Darreus Hsiao and Virgus Hong’s majors belong to the same field?
 Answer: Darreus Hsiao’s major is Dentistry. Dentistry belongs to Medicine. Virgus Hong’s major is Nursing. Nursing belongs to Medicine. Medicine and Medicine are the same.
 The answer is: YES
-M_NR Question: What is the sequence of odd-positioned letters in the first word of Darreus Hsiao’s major name?
+Novel Reasoning Question: What is the sequence of odd-positioned letters in the first word of Darreus Hsiao’s major name?
 Answer: Darreus Hsiao’s major is Dentistry.
@@
 To examine factual hallucinations caused by training with tasks containing new knowledge, we need to discriminate known and unknown knowledge, control their usage during training, and evaluate related hallucinations.
-Since initially the model has no exposure to any knowledge of our synthetic dataset, we prepare the study by continue-pretraining the model with a subset of the knowledge, which becomes known to the model; and keep another subset of the knowledge as unknown.
-By mixing the constructed questions from known and unknown-knowledge in varying proportions, we are able to create situations where different proportion of newly introduced knowledge participates in training.
+Since initially the model has no exposure to any knowledge of our synthetic dataset, we prepare the study by continue pre-training the model with a subset of the knowledge, which becomes known to the model; and keep another subset of the knowledge as unknown.
+By mixing the constructed questions from known and unknown knowledge in varying proportions, we are able to create situations where different proportion of newly introduced knowledge participates in training.
 To evaluate how training leads to hallucinations, we reserve another subset of knowledge as test knowledge.
-The test knowledge are continue-pretrained together with the known knowledge during the preparation, but are kept away from further training.
-Therefore, the difference in performance on test set before and after training indicates the influence of factual hallucinations affected by training.
-In addition, we use the real-world ENTITYQUESTIONS dataset (Sciavolino et al.
-2021) derived from Wikidata (Vrandečić and Krötzsch 2014) (denoted as wiki) as an out-of-distribution (OOD) test set to provide a more robust evaluation.
+The test knowledge are continue pre-trained together with the known knowledge during the preparation, but are kept away from further training.
+Therefore, the difference in performance on test set with and without unknown knowledge in the training data indicates the influence of factual hallucinations induced by training new knowledge.
+In addition, we use the real-world ENTITYQUESTIONS dataset (Sciavolino et al., 2021) derived from Wikidata (Vrandečić and Krötzsch, 2014) (denoted as Wiki) as an out-of-distribution (OOD) test set to provide a more robust evaluation.
 3.3 Models and Setups
-We conduct experiments primarily using the Qwen2.5-1.5B model (Team 2024).
-As supplementary validation, we also perform partial experiments on Qwen3-8B (Team 2025) and Llama3.2-1B (Grattafiori et al.
-2024) to assess generalization across model scales and architectures, with their results provided in Appendix G.
+We conduct experiments primarily using the Qwen2.5-1.5B model (Team, 2024).
+As supplementary validation, we also perform key experiments on Llama3.2-1B (Grattafiori et al., 2024), Qwen3-8B-Base (Team, 2025) and Qwen2.5-32B (Team, 2024) to assess generalization across model scales and architectures, with their results provided in Appendix G.
 As our experiments are conducted on base models, we first apply SFT to endow them with the ability to answer questions in the evaluation sets.
@@
 Given that all test knowledge are known to the model, and the training and testing formats are consistent, there are no cases where the answers are correct but incorrectly formatted.
-We report the standard deviation of accuracy where applicable.
+To ensure the generality of our conclusions, we report the standard deviation of accuracy where applicable.
 4 Hallucination Analysis
 Using the Biography-Reasoning dataset, we conduct a systematic study on factual hallucinations induced by learning different tasks containing various types of new knowledge through SFT.
+We additionally report the impact of learning new knowledge during CPT on hallucinations in Appendix E.
 4.1 Knowledge QA
 In this section, we analyze the impact of training on new knowledge in QA tasks.
-The baseline model is trained on samples constructed from the known knowledge of all four types.
+A baseline model is trained on samples constructed from the known knowledge of all four types.
 We then replace the knowledge of one entire type with unknown samples while keeping the other three types unchanged, resulting in four variant models.
-For each variant, we evaluate performance on three groups of QA test sets: (1) Same-Type QA (STQA): the test set corresponding to the replaced knowledge type; (2) Different-Type QA (DTQA): test sets of the remaining three types; (3) wiki: the real-world QA test set.
+For each variant, we evaluate performance drop compared to the baseline model on three groups of QA test sets: (1) Same-Type QA (STQA): the QA test set whose knowledge type matches the type trained with unknown knowledge; (2) Different-Type QA (DTQA): the QA test sets whose knowledge types differ from the type trained with unknown knowledge; (3) Wiki: the real-world QA test set used as OOD evaluation.
 STQA DTQA Wiki
--56.40 ± 4.28 -1.06 ± 0.30 -9.17 ± 4.17
-Table 2: Average performance degradation (%, mean ± std) of four model variants, measured against the baseline model on different QA test groups. Detailed numerical results are reported in Appendix C.
+-54.26 ($\pm$ 3.30) -2.67 ($\pm$ 1.58) -17.72 ($\pm$ 6.56)
+Table 2: Average performance degradation (%, mean with standard deviation) of four model variants. Detailed numerical results are reported in Appendix C.1.
 Learning new knowledge induces factual hallucinations within the same type, with some spillover effects to other types.
@@
 Training on unknown knowledge leads to substantial performance drops on the STQA test set, reducing the accuracy by more than half.
-We also observe cross-type degradation, as training on one type negatively impacts average performance on others, including the real-world wiki test set containing purely OOD knowledge.
+We also observe cross-type degradation, as training on one type negatively impacts average performance on others, including the real-world Wiki test set containing OOD knowledge.
 This confirms that learning new knowledge can induce hallucinations even on unrelated knowledge.
@@
 We further investigate how varying the proportion of unknown knowledge within a single type influences hallucination tendencies.
-Starting from the fully known-knowledge baseline, we progressively replace 5%, 10%, 20%, 50%, 80%, and 100% of the knowledge in one type with its corresponding unknown knowledge, while still keeping the other three types entirely known.
+Starting from the fully known-knowledge baseline, we progressively replace 5%, 10%, 20%, 50%, 80%, and 100% of the knowledge in one type with unknown knowledge, while still keeping the other three types entirely known.
 Two strategies are considered for handling the remaining known knowledge within the modified type: KeepKnown, where the remaining known instances are retained, and RemoveKnown, where they are excluded from training.
-Figure 2: Performance under two settings with different proportions of unknown knowledge in the same type and wiki test set.
+Figure 2: Performance under two settings with different proportions of unknown knowledge in the same type test set and Wiki test set.
 As shown in Figure 2, the results across the four subplots are mutually corroborative, revealing a consistent pattern: the higher the proportion of unknown knowledge, the more severe the hallucination.
-In KeepKnown, performance on both the in-type test set and the OOD wiki test set degrades gradually at first, followed by a sharp decline as the amount of unknown knowledge increases.
-While RemoveKnown exhibits a more rapid degradation: within the target type, accuracy drops nearly linearly with the ratio of unknown samples; on the wiki set, performance declines sharply at low unknown counts and then plateaus.
-For the same replacement ratio, the key difference between KeepKnown and RemoveKnown lies in whether the type still contains known knowledge.
-We observe that this distinction has a substantial impact on model performance: RemoveKnown consistently underperforms KeepKnown.
-These observations suggest that sparse but fully unknown types are more disruptive than those containing a mixture of known and unknown knowledge, which differs from previous common understanding and poses new challenges for hallucination mitigation.
+In KeepKnown, performance on both the same-type test set and the OOD Wiki test set degrades gradually at first, followed by a sharp decline as the unknown knowledge dominants.
+In contrast, RemoveKnown exhibits a much faster degradation: even at low replacement ratios, the model already shows severe hallucination effects.
+For the same replacement ratio, the key difference between KeepKnown and RemoveKnown lies in whether the knowledge type still contains any known instances.
+We observe that this distinction has a substantial impact on model performance, with RemoveKnown consistently underperforming KeepKnown.
+These observations suggest that sparse but fully unknown knowledge types are more disruptive than those containing a mixture of known and unknown knowledge, which differs from prior common understanding.
+We additionally conduct the same set of experiments in this subsection using the real-world ENTITYQUESTIONS dataset.
+The results are reported in Appendix C.2, and all findings are highly consistent with those obtained on the synthetic data.
 4.2 Knowledge-based Reasoning
@@
 We then replace one reasoning task with instances derived from unknown knowledge and keep all other unchanged, resulting in 12 variant models.
-We investigate how training on a knowledge-based reasoning task with unknown knowledge affects performance across different downstream tasks.
-Specifically, we examine six distinct test groups: (1) Same-Type Same-Reasoning (STSR): the exact reasoning task that trained with unknown knowledge type; (2) Same-Type Different-Reasoning (STDR): different reasoning tasks within the same knowledge type; (3) Different-Type Different-Reasoning (DTDR): all other reasoning tasks with different knowledge; (4) Same-Type QA (STQA): the QA task with the same knowledge type as STSR; (5) Different-Type QA (DTQA): QA tasks with other three knowledge types; and (6) wiki: the real-world QA test set.
-We measure the relative performance change with respect to the baseline and compute the average difference within each of the six task groups.
+We investigate how training on a knowledge-based reasoning task with unknown knowledge affects performance across different downstream tasks, including both knowledge-based reasoning and knowledge QA.
+Specifically, we examine three groups of reasoning tasks: (1) Same-Type Same-Reasoning (STSR): the exact reasoning task that trained with unknown knowledge type; (2) Same-Type Different-Reasoning (STDR): different reasoning tasks within the unknown knowledge type; and (3) Different-Type Different-Reasoning (DTDR): all other reasoning tasks with different knowledge types.
+We also evaluate the model on the knowledge QA test groups defined in Section 4.1, namely STQA, DTQA, and Wiki.
+We measure the relative performance change with respect to the baseline model and compute the average difference within each of the six task groups.
 Results in Figure 3 show that learning reasoning tasks with new knowledge consistently induces performance degradation across all six groups.
-The overall trend aligns with previous findings: the most severe hallucinations occur in STSR, indicating strong intra-task interference, while hallucination on other task groups are relatively minor.
-A notable difference is that among other tasks, QA test sets exhibit even stronger hallucinations than seemingly more related reasoning tasks: STQA, DTQA, and even wiki show greater degradation than STDR and DTDR.
-Within QA, the degradation is more pronounced in STQA than in DTQA, implying that shared factual grounding heightens vulnerability.
-Figure 3: The impact of learning new knowledge in reasoning tasks on the average performance across different groups. Results of each variant model on each dataset are presented in Appendix C.
-5 KnownPatch: Mitigating Hallucination
-In Section 4.1, we find that when all factual knowledge in a certain type is new, even if it constitutes a small proportion of the overall data, it significantly increases the risk of hallucination.
-Prior work typically attempts to mitigate hallucinations by filtering out all data that contain unknown knowledge.
-However, in real-world settings, factual knowledge spans multiple categories and is intertwined with diverse task contexts, making it difficult to separate knowledge from task-related information.
-Identifying which portions of the data involve unknown knowledge is often ambiguous, and discarding entire samples that might contain it can be highly inefficient.
-As a result, perfect filtering is neither feasible nor effective in practice, and the remaining new knowledge can still induce serious hallucinations.
-Even if perfect filtering were achievable, one might expect that transferring the new knowledge to Continued Pre-Training (CPT) would allow the model to first acquire it safely before SFT.
-However, our experiments show that learning new knowledge during the CPT phase can also cause hallucination behaviors similar to the observations in Section 4.1, with detailed results provided in Appendix E.
-To address this issue, we propose KnownPatch, a simple yet effective approach that stabilizes the model by briefly reinforcing known knowledge at the final stage of training.
-5.1 Intuition
-We observe that learning tasks with known knowledge does not introduce instability, whereas training on unknown knowledge can easily disrupt the model’s behavior.
-Based on this observation, KnownPatch places a small portion of known knowledge at the very end of the training sequence.
-The idea is to allow the model, after being inevitably disturbed by new knowledge, to recover a stable state through a brief reinforcement of previously known information.
-This approach is lightweight and easy to apply, as it only requires identifying a small subset of known samples from the SFT corpus.
-While precise separation of all known and unknown knowledge is practically infeasible, extracting a small subset of clearly known data is straightforward (Gekhman et al.
-2024), making KnownPatch simple yet effective in stabilizing the overall training process.
-5.2 Setups
-KnownPatch modifies the training sequence by placing a small subset of SFT data with all known knowledge at the end of the training.
-In our setting, the data trained prior to this subset consists entirely of unknown knowledge across all of the types, to simulate the worst situation.
-We consider two scenarios: (1) the ideal scenario, where the injected known knowledge covers all knowledge types that are involved during the previous training; (2) a more realistic scenario, where the injected known knowledge does not cover all types, and we specifically examine the case where known data from only one type is missing.
-We evaluate KnownPatch across injection ratios of 5%, 10%, and 20% under both scenarios.
-The baseline model is trained on randomly shuffled data, where most knowledge within each type remains unknown; the theoretical upper bound is a model trained solely on known knowledge with the same dataset size.
-To present the observations more clearly in the main text, we average and aggregate certain results, while detailed results are provided in Appendix F.
-In the first scenario, we average the results across the four test types (B, D, M, U) and collectively refer to them as QA, while the OOD wiki results, which are not involved in training, are reported separately.
-In the second scenario, we train four variant models, each corresponding to a different missing type in the injected known knowledge.
-We then average the performance of each variant on its respective missing-type test set and refer to this average as QA, while the averaged wiki performance across the four variants is reported separately.
-5.3 Results
-Figure 4: Averaged performance and attention score changes under different settings.
-The two bar groups on the left represent the performance of KnownPatch in the first scenario, while the bar group on the right corresponds to the second scenario.
-Values here represent the performance change percentage compared to the upper-bound model.
-KnownPatch effectively mitigates factual hallucinations with minimal data injection.
-The left two groups of bars in Figure 4 show the performance of KnownPatch in knowledge QA tasks when injecting 5% and 20% of known data at the end of training.
-When the injected data covers all previous unknown knowledge types, KnownPatch consistently outperforms the baseline with randomly shuffled data across all injection ratios.
-Even with only 5% injected known data, KnownPatch recovers a substantial portion of the performance lost due to hallucinations.
-At 20%, the model’s performance becomes very close to that of the upper-bound model trained entirely on known knowledge.
-Notably, on the OOD wiki test set, performance with just 5% injected known data already approaches the upper bound, and at 20%, it even exceeds the performance of the fully known model, suggesting that late-stage injection of diverse known facts may enhance generalization and stabilize cross-domain knowledge application.
-Even when one knowledge type is missing from the injection data, KnownPatch still substantially mitigates hallucination.
-As shown in the right group of bars in Figure 4, the model achieves significant performance gains even on the uncovered type, despite never being trained on its known facts.
-On the OOD wiki test set, performance also shows a clear improvement over the shuffled baseline and even comparable to the all-known upper-bound.
-This suggests that KnownPatch induces a global stabilization effect that benefits even types not directly represented in the patch.
-KnownPatch also performs well in reasoning tasks.
-In knowledge-based reasoning tasks, when the injected patch data covers all previously unknown knowledge-based reasoning tasks, KnownPatch also achieves consistent performance improvements across all task types with different injection ratios.
-This effect can also transfer to QA tasks that are affected but not directly injected.
-Figure 5 shows the results of 20%, demonstrating the general effectiveness of the KnownPatch method.
-Figure 5: Performance of KnownPatch on reasoning tasks when injecting 20% known data. The values here represent the accuracy percentage of this model compared to the fully known upper-bound model.
-6 Interpretability with Attention Analysis
-In this section, we analyze the mechanisms under new-knowledge-induced factual hallucinations through the attention scores.
-We measure the relative changes in entity attention and task performance in models trained with different data compositions (using the model trained with entirely known data as the baseline).
-We also analyze the attention pattern changes after applying the KnownPatch method.
-6.1 Analysis Setup
-We analyze the changes in attention scores that the model assigns to key entities in the context when applying knowledge.
-In the Biography-Reasoning dataset, the key entity in each question is the person name.
-Therefore, we examine the model’s attention score on the names when generating the first token of the knowledge.
-Prior interpretability studies suggest that the model processes inputs in three stages: understanding, execution, and output, among which language-agnostic knowledge retrieval and abstract reasoning primarily occur in the middle-to-later layers (Wendler et al.
-2024; Zhao et al.
-2024).
-We also examine the model’s average attention to entities across different layers in both reasoning and knowledge QA tasks (detailed in Appendix D), with results shown in Figure 6, demonstrating that this attention is significantly higher in layers 12–24 (out of 28 total layers in Qwen2.5-1.5B), consistent with prior works.
-Therefore, in this study, we compute the average attention scores over layers 12 to 24 for interpretability analysis.
-We measure the relative change in entity attention to assess how training on new knowledge affects the model’s focus on critical entities, by comparing models trained under some new data to the model trained entirely on known knowledge, i.e.
-the upper bound model.
-Figure 6: Attention score on the target name across layers in QA and reasoning training setups.
-The solid curves show the average attention score at each layer, aggregated across all datasets and instances.
+The overall trend aligns with previous findings: the most severe hallucinations occur in STSR, indicating strong intra-task interference.
+Moreover, among other tested tasks, QA test sets exhibit even stronger hallucinations than several seemingly more related reasoning tasks: STQA, DTQA, and even the Wiki test set show greater performance degradation than STDR and DTDR.
+Figure 3: The impact of learning new knowledge in reasoning tasks on the average performance across different groups. Detailed results are presented in Appendix C.1.
+5 Interpretability Analysis
+In this section, we analyze the underlying mechanisms of new-knowledge-induced factual hallucinations by examining the relative changes in attention patterns and hallucination severity.
+Based on this analysis, we propose a simple training intervention, referred to as KnownPatch, which injects a small amount of known knowledge at the later training stage to restore attention patterns and alleviate hallucination behavior.
+Crucially, we also investigate the role of contextual similarity, providing evidence that hallucinations propagate primarily through shared lexical contexts.
+5.1 Attention Analysis Setup
+We measure how learning new knowledge alters the model’s attention to key entities.
+In the Biography-Reasoning dataset, the key entities are person names, so we quantify the model’s attention to the name tokens when generating the first token of the related knowledge.
+Prior interpretability works suggest that knowledge retrieval and reasoning occur primarily in mid-to-late transformer layers (Wendler et al., 2024; Zhao et al., 2024).
+We confirm this trend in our model by examining attention across layers in both QA and reasoning settings (as detailed in Appendix D).
+Figure 4 shows that attention on key entities peaks in layers 12-24 (out of 28 layers), so we average attention over these layers in all subsequent analyses.
+We measure the relative change in entity attention by comparing models trained under new knowledge to the model trained entirely on known knowledge, i.e.
+the baseline model.
+Figure 4: Attention score on the key entity name across layers in QA and reasoning training setups.
+The solid curves show the average attention score aggregated across all instances.
 The shaded regions represent the standard deviation.
-We use this method to analyze (1) the performance degradation of QA tasks under different proportions of unknown knowledge (Figure 2), (2) the impact of learning new knowledge in reasoning tasks (Figure 3), and (3) the effect of applying KnownPatch (Figure 4).
-6.2 Analysis Results and Observations
-Figures 7, 8 and 9 show the changes in entity attention and task performance for models trained on unknown knowledge in various settings.
-Hallucinations emerge alongside a concurrent decline in entity attention and performance, and this coupling effect intensifies in the absence of known knowledge.
+5.2 Correlation between Attention and Hallucination
+STQA DTQA Wiki
+Acc. Attn. Acc. Attn. Acc. Attn.
+w/o attnKL -54.26 ($\pm$ 3.30) -28.02 ($\pm$ 10.08) -2.67 ($\pm$ 1.58) -5.29 ($\pm$ 4.08) -17.72 ($\pm$ 6.56) -7.33 ($\pm$ 3.03)
+w/ attnKL -22.22 ($\pm$ 11.70) -14.18 ($\pm$ 6.09) -5.45 ($\pm$ 3.68) -11.69 ($\pm$ 7.88) -7.13 ($\pm$ 6.57) -10.14 ($\pm$ 1.12)
+Table 3: Results of adding KL loss.
+“Acc.” and “Attn.” denote the averaged performance drop and attention drop on key entities (%, mean with standard deviation), respectively.
+The baseline model is the same as Section 4.1.
+Hallucinations correlate with declines in entity attention.
 Figure 5 presents the interpretability analysis corresponding to Figure 2.
 As the proportion of unknown instances within a knowledge type increases, the model’s attention to key entities gradually declines, accompanied by more severe hallucinations.
-The decline in entity attention under RemoveKnown is faster and leads to a more abrupt performance drop than under KeepKnown, indicating that the absence of known information within a knowledge type accelerates attention decay, which is strongly correlated with performance degradation.
+Compared to KeepKnown, RemoveKnown exhibits a sharper attention decline and performance drop, indicating that the absence of known information accelerates attention decay and exacerbates performance degradation.
+Figure 6 shows the interpretability analysis for the reasoning tasks corresponding to Figure 3, where attention patterns are also correlated with hallucinations.
+Similar experiments are also conducted on the real-world ENTITYQUESTIONS dataset. The results are reported in Appendix C.2.
 Figure 5: Accuracy and attention score changes with different unknown data ratio in certain type.
-Contextual similarity plays a crucial role in how hallucinations propagate across tasks.
-Figure 8 presents the interpretability analysis corresponding to Figure 3.
-We observe that both the performance drop and the reduction in attention to key entities are most pronounced in STSR, followed by the QA tasks.
-Overall, these two metrics exhibit a strong correlation across different task groups.
-Table 3 further shows the contextual similarity between STSR and other task groups.
-On average, QA tasks demonstrate the highest contextual similarity to the reasoning task involving unknown knowledge.
-Owing to their relatively long input contexts, reasoning tasks tend to be less similar to each other, whereas knowledge QA contexts often overlap with substrings of reasoning contexts.
-Taken together, these findings indicate that performance degradation grows with contextual similarity, implying that inter-task hallucinations are likely to spread through shared contextual patterns.
-Consequently, learning new knowledge in reasoning tasks is less likely to induce hallucinations in other reasoning tasks, while QA tasks are more vulnerable to such interference.
-Figure 8: Accuracy and attention score changes compared to the all-known baseline model when learning new knowledge in reasoning tasks.
+Figure 6: Accuracy and attention score changes when learning new knowledge in reasoning tasks.
+[…]
+Constraining attention alleviates factual hallucinations.
+Table 3 presents a comparison with Table 2 after introducing the KL constraint.
+We observe that, on the STQA test sets whose knowledge types are consistent with the unknown knowledge in the training data, as well as on the OOD Wiki test set, hallucinations are substantially alleviated.
+However, we also find that for DTQA and the wiki test set, attention on key entities dropped more even after applying the KL constraint.
+To further investigate this phenomenon, we apply the same KL constraint to train another baseline model (trained entirely on known knowledge).
+This results in an average accuracy drop of 3.67% and a 9.21% reduction in attention to key entities on the test sets, compared to the normally trained baseline model.
+This indicates that training on known knowledge naturally encourages increased attention to key entities, which we believe is beneficial for learning the task.
+Consequently, forcing attention to remain unchanged inevitably suppresses this, leading to a modest reduction in entity attention and a corresponding loss in accuracy.
+Overall, these results suggest that attention shifts during new knowledge training are a key contributing factor to the emergence of hallucinations.
+5.3 KnownPatch: Late-stage Injection of Known Knowledge
+Building on the above analysis, which shows that training on known knowledge promotes increased attention to key entities, we further explore whether injecting a small amount of known knowledge at the final stage of training can help alleviate factual hallucinations induced by learning new knowledge.
+We refer to this training method as KnownPatch.
+The underlying intuition is that exposure to unfamiliar knowledge can disrupt the model’s attention patterns, whereas re-introducing known knowledge encourages the restoration of entity-centric attention and leads to more stable model behavior.
+The training data used before applying the fully known injection patch consist entirely of unknown knowledge across all types, simulating a worst-case scenario.
+We define the injection ratio as the proportion of injected known samples relative to the total training data, and we experiment with injection ratios of 5%, 10%, and 20%.
+The baseline model is trained on known knowledge of the full training data size, serving as an upper bound.
+To control for the effect of training order, we additionally compare against a model trained on the same mixed data after shuffling.
+Our analysis focuses on the relative performance and attention changes of the models trained under the KnownPatch and Shuffled settings compared to the baseline model.
+Figure 7: Performance and attention score changes under Shuffled and KnownPatch (with 20% injection ratio) settings.
+QA represents the average across the four QA test sets, and error bars indicate standard deviations.
+KnownPatch stabilizes the model’s attention patterns and mitigates hallucinations.
+Figure 7 reports the performance drop and the change in entity attention on the QA test sets under a 20% injection ratio.
+Across both in-domain QA and the OOD wiki set, KnownPatch consistently outperforms the shuffled baseline.
+With 20% injection, QA performance approaches the all-known upper bound, and performance on the OOD Wiki set even slightly exceeds it.
+Results under different injection ratios and reasoning tasks are provided in Appendix F, and we find that even with only 5% injection, KnownPatch already yields substantially higher performance than the Shuffled setting on both QA and Wiki test sets.
+These gains are accompanied by increased attention to key entities in the questions, suggesting that KnownPatch restores entity-centric attention patterns disrupted by training on unknown knowledge and yields a robust mitigation effect that generalizes to OOD data.
+5.4 Mechanism of Hallucination Propagation
+We further aim to investigate how new-knowledge-induced factual hallucinations propagate across different tasks.
 STSR STDR DTDR STQA DTQA wiki
-1.0000 0.6164 0.5896 0.7312 0.6982 0.7199
-Table 3: Averaged contextual similarity between STSR and other test groups, where each value denotes the mean similarity between the STSR context and all task contexts within the corresponding group.
-Similarity between A and B is computed as the proportion of tokens in B appearing in A.
-KnownPatch stabilizes the model’s attention patterns and mitigates hallucinations.
-Figure 9 shows the performance drop and changes in attention on QA test sets relative to the all-known baseline model, providing an interpretability analysis corresponding to the right group of bars in Figure 4.
-It can be observed that KnownPatch can effectively restore the model’s focus on key entities in the questions, thereby significantly alleviating the hallucination effects caused by learning new knowledge.
-Figure 9: Performance and attention score changes when learning new knowledge in QA tasks, and after applying KnownPatch (with 20% known data).
-QA represents the average across the four QA test sets, and error bars indicate standard deviations.
+1.00 0.62 0.59 0.73 0.70 0.72
+Table 4: Averaged token-level overlap between STSR and other test groups.
+Each value represents the mean token overlap ratio, defined as the proportion of tokens in a task context that also appear in the STSR context, averaged over all tasks in the corresponding group.
+Since attention weights form a normalized distribution that sums to one across all input tokens, a reduction in attention to key entities implies a corresponding shift of attention toward the remaining contextual tokens.
+Figure 3 shows that learning new knowledge in reasoning tasks has a more substantial impact on QA test sets than on other reasoning test sets.
+This is consistent with the lexical similarity (measured by token-level overlap) analysis in Table 4: QA test sets are generally more lexically similar to the STSR tasks that trained with unknown knowledge, and this higher lexical similarity aligns with the observed trend of stronger hallucination propagation.
+This can be attributed to the fact that reasoning tasks typically involve relatively long and diverse input contexts, making them lexically less similar to one another, whereas many knowledge QA contexts share substrings with reasoning trajectories.
+Lexical Same Similar Different
+Semantic Same Same Diff. Same Diff.
+Token Overlap 1.00 0.97 0.89 0.62 0.52
+Hallucination 25.57 18.80 16.61 6.65 5.90
+Table 5: Averaged contextual similarity (token overlap) and performance drop (hallucination, %).
+To disentangle the effects of lexical similarity and semantic similarity in contexts on hallucination propagation, we construct extended variants of the reasoning tasks.
+Specifically, for each of the 12 reasoning tasks in Biography-Reasoning, we construct four variant tasks under the same knowledge type, defined by lexical similarity (similar vs.
+different) and semantic similarity (same vs.
+different).
+An example for each variant task is provided in Appendix A.
+We then examine, across the four variant tasks, the extent of performance degradation when the Origin task is trained on unknown knowledge compared to training on known knowledge.
+Results in Table 5 show that learning new knowledge induces stronger hallucinations on tasks with higher lexical similarity.
+This confirms that hallucinations propagate primarily through lexical similarity rather than semantic relatedness.
 6 Conclusion
-In this work, we present a systematic study on hallucinations caused by learning new knowledge in LLMs, examining their behavior across knowledge types and task types.
-Our experiments reveal that even a small number of fully unknown facts can trigger severe hallucinations, not only in the target task but also in other unrelated knowledge QA tasks, indicating a non-trivial inter-task impact.
-To address this hallucination, we propose KnownPatch, a simple but practical mitigation strategy that injects a small proportion of known knowledge samples during the final phase of fine-tuning.
-Without extensive data, KnownPatch effectively reduces hallucination across both QA and reasoning tasks even under partial type coverage.
-Through further analysis, we find that learning new knowledge shifts the model’s focus from key entities to other context, promoting erroneous binding and hallucination propagation.
-And this altered attention pattern can propagate along similar contexts.
-KnownPatch counteracts this by restoring entity-centric attention patterns.
+In this work, we present a systematic study on new-knowledge-induced factual hallucinations in LLMs, examining their behavior across knowledge types and task types.
+Our experiments reveal that even a small number of fully unknown facts can trigger severe hallucinations and can propagate to other tasks.
+Our analysis reveals that this behavior is closely associated with shifts in attention: learning new knowledge reduces attention to key entities, whereas training on known knowledge reinforces entity-centric attention.
+Motivated by this observation, we explore a simple training intervention, KnownPatch, which injects a small amount of known knowledge at the late stage of training to restore attention patterns and mitigate hallucinations.
+Finally, we show that the extent of hallucination propagation increases with lexical similarity between contexts, rather than semantic relatedness, highlighting contextual similarity as a key driver of cross-task hallucination transfer.
+Limitations
+While our work offers a in-depth analysis of new-knowledge-induced factual hallucinations, there are several boundaries to our current study.
+Regarding the synthetic data:
+To strictly disentangle new knowledge from facts already internalized during pre-training, we constructed the synthetic Biography-Reasoning dataset.
+While this controlled environment was necessary, it may not fully capture the noise and semantic complexity of naturally occurring text.
+We partially mitigated this by validating our findings on the real-world ENTITYQUESTIONS dataset, yet exploring more complex linguistic structures remains a direction for future work.
+Regarding model scale:
+Our analysis primarily relies on open-source models (e.g., Qwen2.5-1.5B) to enable detailed and fine-grained investigation.
+While we verify the key findings on larger variants (e.g., Qwen3-8B, Qwen2.5-32B) in the appendix, we acknowledge that extending our experiments to extremely large-scale models (e.g., 70B+ or even larger models) is left for future work due to computational constraints.
 Appendix A Dataset Details
@@
 The ENTITYQUESTIONS dataset from Wikidata is divided into multiple subsets, such as P17, P20, etc., with each subset containing questions of the same format.
-For example, an instance from P17 is “Which country is Juniper Bank located in?”, and an instance from P20 is “Where did Connee Boswell die?”.
+For example, an instance from P17 is “Which country is Juniper Bank located in?", and an instance from P20 is “Where did Connee Boswell die?".
 Based on Gekhman et al.
@@
 The first name and last name of each individual are selected from separate pools and are ensured to be unique.
-For first names, we use 3,000 English names with an equal split between male and female names (this affects the use of gendered pronouns in reasoning tasks).
-For last names, we select 250 Chinese surnames , which are then randomly paired with the first names in a balanced manner.
+For first names, we use 3,000 English names from the UCI Machine Learning Repository dataset [^2] https://archive.ics.uci.edu/dataset/591/gender+by+name, which is under CC BY 4.0 license and could be used for any purpose., with an equal split between male and female names (this affects the use of gendered pronouns in reasoning tasks).
+For last names, we select 250 Chinese surnames from a GitHub repository [^3] https://github.com/smashew/NameDatabases/, which is under ‘The Unlicense’ that allows anyone to use for free.
+, which are then randomly paired with the first names in a balanced manner.
 This random combination of English first names and Chinese last names is designed to generate synthetic individuals that minimize overlap with real-world knowledge already known to language models.
@@
 There are also 50 majors, categorized into 10 broad fields (5 majors per field), e.g., Computer Science → Engineering.
+We refer to the four attributes Birth, Death, Major and University as B, D, M and U, respectively.
 CPT Data
@@
 To ensure that the model reliably retains them, we also rephrase each auxiliary fact 50 times and include them in the CPT data.
-All auxiliary knowledge is provided in Tables 5 and 6.
+All auxiliary knowledge is provided in Tables 8 and 9.
 Reasoning Tasks
-For each attribute of a synthetic individual, we construct three types of reasoning questions. In Table 4, we provide an example for each category of QA and reasoning questions in the dataset.
+For each attribute of a synthetic individual, we construct three types of reasoning questions.
+We refer to Single Reasoning, Comparative Reasoning, and Novel Reasoning as SR, CR, and NR, respectively.
+In Table 6, we provide an example for each category of QA and reasoning questions in the dataset.
 Each CR task involves two attributes: the primary attribute of interest and another randomly selected one.
 For major- and university-related CR tasks, which take a binary (Yes/No) form, we further constrain the sampling process to maintain an approximately balanced ratio of positive and negative instances (50% each).
-During the SFT stage for reasoning tasks, we split the known individual set into two subsets: 80% are used to generate reasoning questions, and the remaining 20% are used for QA tasks.
-This allows the model to be exposed to both task types, ensuring a more reliable evaluation across reasoning and QA.
+To investigate the effect of context similarity on hallucination propagation, we construct variant tasks for each reasoning problem under the same knowledge type: (a) Same Meaning, Similar Lexical (SMSL); (b) Same Meaning, Different Lexical (SMDL); and (c) Different Meaning, Similar Lexical (DMSL).
+Table 7 presents representative examples of these variants.
+For the Different Meaning, Different Lexical setting, we directly adopt the STDR tasks defined in Section 4.2.
 Category Example
@@
 Table 6: Examples of each QA and reasoning tasks in Biography-Reasoning.
+B, D, M, and U denote birth year, death year, major, and university, respectively; SR, CR, and NR denote Single Reasoning, Comparative Reasoning, and Novel Reasoning, respectively.
+Category Example
+B_SR SMSL: Is Hakam Cheng’s birth year an odd number?
+SMDL: Can Hakam Cheng’s year of birth, when considered as an integer, be classified under the category of odd values?
+DMSL: Is the number of Hakam Cheng’s birth year an even number?
+B_CR SMSL: How many years is the birth year between Hakam Cheng and Graicyn Xian apart?
+SMDL: By what number of years are Hakam Cheng and Graicyn Xian separated in terms of their birth years?
+DMSL: How many months apart is the birth year between Hakam Cheng and Graicyn Xian?
+B_NR SMSL: What’s the MScore of Hakam Cheng’s birth year?
+SMDL: Determine the specific MScore value attributed to the calendar year during which Hakam Cheng was born.
+DMSL: What is the AScore of Hakam Cheng’s birth year?
+D_SR SMSL: What is the 10th anniversary of Hakam Cheng’s death year?
+SMDL: After 10 years from Hakam Cheng passed away, what year does that correspond to?
+DMSL: What is the year before the 10th anniversary of Hakam Cheng’s death?
+D_CR SMSL: Hakam Cheng or Graicyn Xian, who died first?
+SMDL: Identify which individual—Hakam Cheng or Graicyn Xian—died at an earlier date.
+DMSL: Who died later, Hakam Cheng or Graicyn Xian?
+D_NR SMSL: What’s the AScore of Hakam Cheng’s death year?
+SMDL: Which AScore value corresponds to the calendar year in which Hakam Cheng died?
+DMSL: What is the MScore of Hakam Cheng’s death year?
+M_SR SMSL: Which field does Hakam Cheng’s major belong to?
+SMDL: To which academic discipline can the major pursued by Hakam Cheng be categorized?
+DMSL: What is the first letter of the field that Hakam Cheng’s major belongs to?
+M_CR SMSL: Do Hakam Cheng’s and Graicyn Xian’s majors belong to the same field?
+SMDL: Are Hakam Cheng and Graicyn Xian’s areas of academic specialization considered part of the same disciplinary category?
+DMSL: Do Hakam Cheng and Graicyn Xian’s majors belong to different field?
+M_NR SMSL: What is the sequence of odd-positioned letters within the first word of Hakam Cheng’s major name?
+SMDL: From the initial word of Hakam Cheng’s major, extract the characters occupying positions with odd indices and present them in order.
+DMSL: What is the sequence of even-positioned letters in the first word of Hakam Cheng’s major name?
+U_SR SMSL: Hakam Cheng attend university in which country?
+SMDL: Identify the nation within whose borders Hakam Cheng pursued university-level studies.
+DMSL: What is the last letter of the country in which did Hakam Cheng attended university?
+U_CR SMSL: Hakam Cheng and Graicyn Xian are college alumni?
+SMDL: Have Hakam Cheng and Graicyn Xian both completed their higher education at the same university?
+DMSL: Are Hakam Cheng and Graicyn Xian not college alumni?
+U_NR SMSL: What’s the sequence of the first and last letters of each word in Hakam Cheng’s university name?
+SMDL: Provide the ordered sequence formed by the initial and final characters of every word appearing in Hakam Cheng’s university title.
+DMSL: What is the sequence of the last and first letters of each word in Hakam Cheng’s university name?
+Table 7: Examples of the variant tasks for reasoning tasks.
 Appendix B Training Details
@@
 We also did experiments of training 1 or 5 epochs, the results are presented in Appendix H.
-All experiments are conducted on up to 4 NVIDIA A6000 GPUs.
-The CPT stage is performed using LLaMA-Factory (Zheng et al.
-2024).
+In the experiment that uses an auxiliary KL loss in Section 5.2, due to this additional constraint, training becomes more difficult, so we set the number of training epochs to at most ten and apply early stopping when the training accuracy exceeds 95%.
+All other training setups are the same as above.
+Most of the experiments are conducted on up to four NVIDIA A6000 GPUs for models with fewer than 8B parameters.
+For training models with 8B parameters or more, up to eight NVIDIA H20 GPUs are used.
+The CPT stage is performed using LLaMA-Factory (Zheng et al., 2024).
 Field Major
@@
 Table 11: Impact on other QA test sets when training new knowledge in reasoning tasks.
-Appendix C Detailed Results of Main Text
+Appendix C More Results for Main Text
+C.1 Detailed Results for Main Text
 In this section, we detail the test results of each model on each dataset as shown in Table 2 and Figure 3 in the main text.
@@
 Model B_QA D_QA M_QA U_QA wiki
-$\text{All}_{\text{k}}$ 0.549 0.609 0.546 0.464 0.199
-$\text{B}_{\text{unk}}$ 0.200 0.591 0.555 0.466 0.195
-$\text{D}_{\text{unk}}$ 0.539 0.225 0.531 0.464 0.188
-$\text{M}_{\text{unk}}$ 0.545 0.596 0.255 0.456 0.183
-$\text{U}_{\text{unk}}$ 0.531 0.606 0.546 0.252 0.157
+$\text{All}_{\text{k}}$ 0.549 0.650 0.519 0.442 0.199
+$\text{B}_{\text{unk}}$ 0.252 0.618 0.513 0.430 0.179
+$\text{D}_{\text{unk}}$ 0.521 0.322 0.511 0.426 0.168
+$\text{M}_{\text{unk}}$ 0.539 0.627 0.244 0.445 0.165
+$\text{U}_{\text{unk}}$ 0.528 0.635 0.508 0.179 0.143
 Table 12: Hallucination induced by SFT on different unknown knowledge types.
+C.2 Results on Real-World Dataset
+To examine whether our findings hold on real-world data, we additionally conduct experiments on the ENTITYQUESTIONS dataset (Sciavolino et al., 2021).
+A key difference between real-world data and our synthetic dataset lies in whether the corresponding knowledge has already been observed by the model during its initial pre-training stage.
+Following the procedure described in Appendix A.1, we categorize knowledge for the Qwen2.5-1.5B model into four groups: HighlyKnown, MaybeKnown, WeaklyKnown, and Unknown.
+We select four knowledge types that contain a relatively large proportion of HighlyKnown and Unknown instances: P36, P106, P159, and P407, and use them as training knowledge types.
+Following the experimental setups in Section 4.1, we sample 500 instances for each knowledge type.
+A model trained on the fully known dataset serves as the baseline model, while variant models are obtained by replacing one knowledge type with unknown instances.
+We then evaluate these models on the corresponding test sets of the four knowledge types, as well as on the same Wiki test set used in the main text.
+The results are shown in Table 13, which shows exactly the same trends as Table 2.
+STQA DTQA wiki
+-36.94 ($\pm$ 11.30) -3.29 ($\pm$ 3.35) -6.86 ($\pm$ 2.57)
+Table 13: Average performance degradation (%, mean ± std) of four model variants.
+We conduct experiments to examine how the proportion of unknown knowledge affects the severity of hallucinations. The results are shown in Figure 8, which shows similar trends as Figure 2.
+Figure 8: Performance with different proportions of unknown knowledge in the same type and Wiki test set.
+Following the experiments in Section 5.2, in Figure 9 we also show that the attention to key entities is correlated to performance drop when unknown knowledge proportion increases.
+Figure 9: Accuracy and attention score changes with different unknown data ratio in certain type.
+Figure 10 presents the results of KnownPatch (with injection ratio 20%) on real-world data, along with its attention analysis.
+Figure 10: Performance and attention score changes under Shuffled and KnownPatch (with 20% injection ratio) settings.
+QA represents the average across the four QA test sets, and error bars indicate standard deviations.
 Appendix D Attention Layer Selection
@@
 If the model answers correctly in at least one case, it is classified as Known; if all answers are incorrect, it is classified as Unknown.
-This is because the selection and order of few shots can significantly affect the model’s performance (Lu et al.
-2022; Zhao et al.
-2021), and we need to rule out this influence.
-The results are shown in Tables 10, 11, 12 and 13.
+This is because the selection and order of few shots can significantly affect the model’s performance (Lu et al., 2022; Zhao et al., 2021), and we need to rule out this influence.
+The results are shown in Tables 14, 15, 16 and 17.
 Among all the results, except for Table 14, there are quite serious hallucination phenomena.
@@
 Appendix F Supplementary Results of KnownPatch
-Figure 10 is the result of KnownPatch on QA tasks with different injection ratios, in addition to Figure 4 in the main text; Figures 11 and 12 are results of KnownPatch on reasoning tasks with injection ratios of 5% and 10%, in addition to Figure 5 (results of injection ratio 5%) in the main text; Figures 13, 14 and 15 are detailed results of KnownPatch when one knowledge type is missed, with injection ratios 20%, 10% and 5%, respectively.
+Figure 11 is the result of KnownPatch on QA tasks with different injection ratios; Figures 12, 13 and 14 are results of KnownPatch on reasoning tasks with injection ratios of 5%, 10% and 20%; Figures 15 and 16 are interpretability results of injecting 10% and 5% known data in KnownPatch.
 Figure 11: Performance of KnownPatch on QA task when injecting 5% (upper), 10% (middle) and 20% (lower) known data. All experiments trained for 3 epoch.
@@
 Figure 13: KnownPatch on reasoning tasks with 10% injection ratio. All experiments trained for 3 epoch.
-Figure 13: KnownPatch (missing one knowledge type) on QA tasks with an injection ratio of 20%. All experiments trained for 3 epoch.
-Figure 14: KnownPatch (missing one knowledge type) on QA tasks with an injection ratio of 10%. All experiments trained for 3 epoch.
-Figure 15: KnownPatch (missing one knowledge type) on QA tasks with an injection ratio of 5%. All experiments trained for 3 epoch.
+Figure 14: KnownPatch on reasoning tasks with 20% injection ratio. All experiments trained for 3 epoch.
+Figure 15: Performance and attention score changes when learning new knowledge in QA tasks, and after applying KnownPatch (with 10% known data).
+QA represents the average across the four QA test sets, and error bars indicate standard deviations.
+Figure 16: Performance and attention score changes when learning new knowledge in QA tasks, and after applying KnownPatch (with 5% known data).
+QA represents the average across the four QA test sets, and error bars indicate standard deviations.
 Appendix G Results on Different Models
@@
 Due to resource limitations, larger-scale training was infeasible.
-For larger models, prior work (Allen-Zhu and Li 2025) shows they retain factual knowledge better.
+For larger models, prior work (Allen-Zhu and Li, 2025) shows they retain factual knowledge better.
 G.1 Results on Llama-3.2-1B
@@
 Table 18 (similar to Table 2) provides the hallucination results in QA tasks when learning new knowledge; Figure 17 (similar to Figure 3) shows the impact of new knowledge in reasoning tasks on different groups.
-Model B_QA D_QA M_QA U_QA wiki
-$\text{All}_{\text{k}}$ 0.863 0.855 0.766 0.668 0.155
-$\text{B}_{\text{unk}}$ 0.332 0.837 0.761 0.657 0.116
-$\text{D}_{\text{unk}}$ 0.836 0.359 0.767 0.664 0.145
-$\text{M}_{\text{unk}}$ 0.845 0.848 0.478 0.664 0.152
-$\text{U}_{\text{unk}}$ 0.855 0.860 0.761 0.387 0.141
+STQA DTQA Wiki
+-49.80 ($\pm$ 11.74) -1.04 ($\pm$ 1.03) -10.65 ($\pm$ 10.11)
 Table 18: Llama-3.2-1B model’s hallucination induced by training on different unknown knowledge types in QA tasks.
@@
 QA represents the average across the four QA test sets, and error bars indicate standard deviations.
-G.2 Results on Qwen3-8B
+G.2 Results on Qwen3-8B-Base
 Due to the large scale of model parameters, our training setting differ from the default one. We only fine-tune for 1 epoch with a learning rate 5e-6 in all the SFT experiments.
 Table 19 (similar to Table 2) provides the hallucination results in QA tasks when learning new knowledge; Figure 20 (similar to Figure 3) shows the impact of new knowledge in reasoning tasks on different groups.
-Model B_QA D_QA M_QA U_QA wiki
-$\text{All}_{\text{k}}$ 0.870 0.864 0.867 0.709 0.301
-$\text{B}_{\text{unk}}$ 0.157 0.850 0.855 0.709 0.253
-$\text{D}_{\text{unk}}$ 0.846 0.165 0.847 0.676 0.277
-$\text{M}_{\text{unk}}$ 0.846 0.855 0.167 0.680 0.285
-$\text{U}_{\text{unk}}$ 0.833 0.852 0.852 0.149 0.229
-Table 15: Qwen3-8B model’s hallucination induced by training on different unknown knowledge types in QA tasks.
-Figure 19: The impact of learning new knowledge in reasoning tasks on the average performance of different groups (on the Qwen3-8B model).
-We perform the same analysis as Section 6 and Appendix D on the Qwen3-8B model.
+STQA DTQA Wiki
+-80.64 ($\pm$ 1.23) -2.33 ($\pm$ 1.42) -13.29 ($\pm$ 8.40)
+Table 19: Qwen3-8B-Base model’s hallucination induced by training on different unknown knowledge types in QA tasks.
+Figure 20: The impact of learning new knowledge in reasoning tasks on the average performance of different groups (on the Qwen3-8B-Base model).
+We perform the same analysis as Section 5 and Appendix D on the Qwen3-8B-Base model.
 Based on the results of Figure 21 (similar to Figure 4), we chose its layers 9-27 for further interpretability analysis, and Figure 22 (similar to Figure 7) shows the results.
-Figure 20: Qwen3-8B model’s attention score on the target name across layers in QA and reasoning training setups.
+Figure 21: Qwen3-8B-Base model’s attention score on the target name across layers in QA and reasoning training setups.
 The solid curves show the average attention score at each layer, aggregated across all datasets and instances.
 The shaded regions represent the standard deviation.
-Figure 21: Qwen3-8B model’s performance and attention score changes when learning new knowledge in QA tasks, and after applying KnownPatch (with 20% known data).
+Figure 22: Qwen3-8B-Base model’s performance and attention score changes when learning new knowledge in QA tasks, and after applying KnownPatch (with 20% known data).
+QA represents the average across the four QA test sets, and error bars indicate standard deviations.
+G.3 Results on Qwen2.5-32B
+Due to the large scale of model parameters, our training setting differ from the default one. We only fine-tune for 1 epoch with a learning rate 5e-6 in all the SFT experiments.
+Table 20 (similar to Table 2) provides the hallucination results in QA tasks when learning new knowledge; Figure 23 (similar to Figure 3) shows the impact of new knowledge in reasoning tasks on different groups.
+STQA DTQA Wiki
+-74.00 ($\pm$ 5.29) -2.59 ($\pm$ 3.57) -9.69 ($\pm$ 7.88)
+Table 20: Qwen2.5-32B model’s hallucination induced by training on different unknown knowledge types in QA tasks.
+Figure 23: The impact of learning new knowledge in reasoning tasks on the average performance of different groups (on the Qwen2.5-32B model).
+We perform the same analysis as Section 5 and Appendix D on the Qwen2.5-32B model.
+Based on the results of Figure 24 (similar to Figure 4), we chose its layers 25-55 for further interpretability analysis, and Figure 25 (similar to Figure 7) shows the results.
+Figure 24: Qwen2.5-32B model’s attention score on the target name across layers in QA and reasoning training setups.
+The solid curves show the average attention score at each layer, aggregated across all datasets and instances.
+The shaded regions represent the standard deviation.
+Figure 25: Qwen2.5-32B model’s performance and attention score changes when learning new knowledge in QA tasks, and after applying KnownPatch (with 20% known data).
 QA represents the average across the four QA test sets, and error bars indicate standard deviations.
@@
 H.1 1 Epoch
-Table 16 (similar to Table 2) provides the hallucination results in QA tasks when learning new knowledge; Figure 22 (similar to Figure 2) shows the performance after learning different proportions of unknown knowledge; Figure 23 (similar to Figure 3) shows the impact of new knowledge in reasoning tasks on different groups; Figure 24 (similar to Figure 10) reports the performance of KnownPatch on QA tasks when injecting 20% known data; Figure 25 (similar to Figure 5) reports performance of KnownPatch on reasoning tasks when injecting 20% known data; Figure 26 (similar to Figure 13) reports performance of KnownPatch when one knowledge type is missing with 20% injection ratio; Figure 27 (similar to Figure 8) reports the accuracy and attention score changes when learning new knowledge in reasoning tasks; Figure 28 (similar to Figure 7) reports the accuracy and attention score changes after learning different proportions of unknown knowledge; Figure 29 (similar to Figure 9) reports the performance and attention score changes before and after applying KnownPatch.
-Model B_QA D_QA M_QA U_QA wiki
-$\text{All}_{\text{k}}$ 0.548 0.568 0.475 0.458 0.198
-$\text{B}_{\text{unk}}$ 0.198 0.527 0.481 0.448 0.199
-$\text{D}_{\text{unk}}$ 0.518 0.280 0.471 0.442 0.190
-$\text{M}_{\text{unk}}$ 0.517 0.576 0.193 0.423 0.173
-$\text{U}_{\text{unk}}$ 0.529 0.594 0.467 0.304 0.174
+Table 21 (similar to Table 2) provides the hallucination results in QA tasks when learning new knowledge; Figure 26 (similar to Figure 2) shows the performance after learning different proportions of unknown knowledge; Figure 27 (similar to Figure 3) shows the impact of new knowledge in reasoning tasks on different groups; Figure 28 (similar to Figure 14) reports performance of KnownPatch on reasoning tasks when injecting 20% known data; Figure 30 (similar to Figure 6) reports the accuracy and attention score changes when learning new knowledge in reasoning tasks; Figure 31 (similar to Figure 5) reports the accuracy and attention score changes after learning different proportions of unknown knowledge; Figure 32 (similar to Figure 7) reports the performance and attention score changes before and after applying KnownPatch.
+STQA DTQA Wiki
+-51.89 ($\pm$ 13.35) -2.53 ($\pm$ 3.72) -7.07 ($\pm$ 6.40)
 Table 21: Hallucination induced by training on different unknown knowledge types in QA tasks. All experiments trained for 1 epoch.
@@
 Figure 27: The impact of learning new knowledge in reasoning tasks on the average performance of different groups. All experiments trained for 1 epoch.
-Figure 24: Performance of KnownPatch on QA task when injecting 20% known data. All experiments trained for 1 epoch.
 Figure 28: Performance of KnownPatch on reasoning task when injecting 20% known data.
@@
 H.2 5 Epochs
-Table 17 (similar to Table 2) provides the hallucination results in QA tasks when learning new knowledge; Figure 30 (similar to Figure 2) shows the performance after learning different proportions of unknown knowledge; Figure 31 (similar to Figure 3) shows the impact of new knowledge in reasoning tasks on different groups; Figure 32 (similar to Figure 10) reports the performance of KnownPatch on QA task with 20% injection ratio; Figure 33 (similar to Figure 5) reports performance of KnownPatch when injecting 20% known data; Figure 34 (similar to Figure 13) reports performance of KnownPatch when one knowledge type is missing when injecting 20% known data; Figure 35 (similar to Figure 8) reports the accuracy and attention score changes when learning new knowledge in reasoning tasks; Figure 36 (similar to Figure 7) reports the accuracy and attention score changes after learning different proportions of unknown knowledge; Figure 37 (similar to Figure 9) reports the performance and attention score changes before and after applying KnownPatch.
-Model B_QA D_QA M_QA U_QA wiki
-$\text{All}_{\text{k}}$ 0.529 0.600 0.552 0.447 0.200
-$\text{B}_{\text{unk}}$ 0.234 0.573 0.550 0.434 0.194
-$\text{D}_{\text{unk}}$ 0.505 0.289 0.533 0.435 0.181
-$\text{M}_{\text{unk}}$ 0.523 0.606 0.215 0.432 0.171
-$\text{U}_{\text{unk}}$ 0.523 0.607 0.554 0.245 0.144
+Table 22 (similar to Table 2) provides the hallucination results in QA tasks when learning new knowledge; Figure 33 (similar to Figure 2) shows the performance after learning different proportions of unknown knowledge; Figure 34 (similar to Figure 3) shows the impact of new knowledge in reasoning tasks on different groups; Figure 35 (similar to Figure 14) reports performance of KnownPatch when injecting 20% known data; Figure 37 (similar to Figure 6) reports the accuracy and attention score changes when learning new knowledge in reasoning tasks; Figure 38 (similar to Figure 5) reports the accuracy and attention score changes after learning different proportions of unknown knowledge; Figure 39 (similar to Figure 7) reports the performance and attention score changes before and after applying KnownPatch.
+STQA DTQA Wiki
+-53.46 ($\pm$ 6.68) -1.79 ($\pm$ 2.05) -13.75 ($\pm$ 10.60)
 Table 22: Hallucination induced by training on different unknown knowledge types in QA tasks. All experiments trained for 5 epoch.
@@
 Figure 34: The impact of learning new knowledge in reasoning tasks on the average performance of different groups. All experiments trained for 5 epoch.
-Figure 32: Performance of KnownPatch on QA task when injecting 5% (upper), 10% (middle) and 20% (lower) known data. All experiments trained for 5 epoch.
 Figure 35: Performance of KnownPatch on reasoning task when injecting 20% known data.
@@
 H.3 20 Epochs
-Table 18 (similar to Table 2) provides the hallucination results in QA tasks when learning new knowledge; Figure 38 (similar to Figure 2) shows the performance after learning different proportions of unknown knowledge; Figure 39 (similar to Figure 3) shows the impact of new knowledge in reasoning tasks on different groups; Figure 40 (similar to Figure 10) reports the performance of KnownPatch on QA task when injecting 20% known data; Figure 41 (similar to Figure 5) reports performance of KnownPatch when injecting 20% known data; Figure 42 (similar to Figure 13) reports performance of KnownPatch when one knowledge type is missing when injecting 20% known data; Figure 43 (similar to Figure 8) reports the accuracy and attention score changes when learning new knowledge in reasoning tasks; Figure 44 (similar to Figure 7) reports the accuracy and attention score changes after learning different proportions of unknown knowledge; Figure 45 (similar to Figure 9) reports the performance and attention score changes before and after applying KnownPatch.
-Model B_QA D_QA M_QA U_QA wiki
-$\text{All}_{\text{k}}$ 0.498 0.611 0.540 0.371 0.186
-$\text{B}_{\text{unk}}$ 0.183 0.503 0.516 0.380 0.176
-$\text{D}_{\text{unk}}$ 0.443 0.231 0.499 0.352 0.159
-$\text{M}_{\text{unk}}$ 0.469 0.544 0.343 0.360 0.178
-$\text{U}_{\text{unk}}$ 0.452 0.549 0.485 0.252 0.171
+Table 23 (similar to Table 2) provides the hallucination results in QA tasks when learning new knowledge; Figure 40 (similar to Figure 2) shows the performance after learning different proportions of unknown knowledge; Figure 41 (similar to Figure 3) shows the impact of new knowledge in reasoning tasks on different groups; Figure 42 (similar to Figure 14) reports performance of KnownPatch when injecting 20% known data; Figure 44 (similar to Figure 6) reports the accuracy and attention score changes when learning new knowledge in reasoning tasks; Figure 45 (similar to Figure 5) reports the accuracy and attention score changes after learning different proportions of unknown knowledge; Figure 46 (similar to Figure 7) reports the performance and attention score changes before and after applying KnownPatch.
+STQA DTQA Wiki
+-48.50 ($\pm$ 16.53) -7.73 ($\pm$ 5.06) -8.06 ($\pm$ 4.58)
 Table 23: Hallucination induced by training on different unknown knowledge types in QA tasks. All experiments trained for 20 epoch.
@@
 Figure 41: The impact of learning new knowledge in reasoning tasks on the average performance of different groups. All experiments trained for 20 epoch.
-Figure 40: Performance of KnownPatch on QA task when injecting 5% (upper), 10% (middle) and 20% (lower) known data. All experiments trained for 20 epoch.
 Figure 42: Performance of KnownPatch on reasoning task when injecting 20% known data.